### v1

# Intrinsic training dynamics of deep neural networks

Abstract
A fundamental challenge in the theory of deep learning is to understand whether gradient-based training in high-dimensional parameter spaces can be captured by simpler, lower-dimensional structures, leading to so-called implicit bias. As a stepping stone, we study when a gradient flow on a high-dimensional variable $\theta$ implies an intrinsic gradient flow on a lower-dimensional variable $z=\phi(\theta)$, for an architecture-related function $\phi$. We express a so-called intrinsic dynamic property and show how it is related to the study of conservation laws associated with the factorization $\phi$. This leads to a simple criterion based on the inclusion of kernels of linear maps which yields a necessary condition for this property to hold. We then apply our theory to general ReLU networks of arbitrary depth and show that, for any initialization, it is possible to rewrite the flow as an intrinsic dynamic in lower dimension that depends only on $z$ and the initialization, when $\phi$ is the so-called path-lifting. In the case of linear networks with $\phi$ the product of weight matrices, so-called balanced initializations are also known to enable such a dimensionality reduction; we generalize this result to a broader class of relaxed balanced initializations, showing that, in certain configurations, these are the only initializations that ensure the intrinsic dynamic property. Finally, for the linear neural ODE associated to the limit of infinitely deep linear networks, with relaxed balanced initialization, we explicitly express the corresponding intrinsic dynamics.
1 Introduction
A central question in deep learning theory is how the complexity of high-dimensional gradient-based training can give rise to simpler, lower-dimensional dynamics. In this work, we explore when the gradient flow on high-dimensional parameters $\theta$ naturally induces a gradient flow on a lower-dimensional variable $z=\phi(\theta)$, where $\phi$ captures structural aspects of the model.
Intrinsic lower dimensional flow
The study of high-dimensional optimization flows arising in the training of neural networks often benefits from the identification of lower-dimensional intrinsic dynamics. Specifically, it is of considerable interest to rewrite a high-dimensional parameter flow $\theta(t)$ in terms of an lower-dimensional representation $z(t)=\phi(\theta(t))$, using a suitable architecture-related reparametrization $\phi$. When the reduced flow on $z(t)$ admits a mirror flow representation, one gains access to powerful analytical tools rooted in convex optimization theory, allowing explicit characterization of the implicit bias induced by gradient-based optimization algorithms. In particular, prior research has successfully leveraged mirror flow formulations to rigorously demonstrate implicit regularization effects, such as sparsity in scalar linear neural networks and two-layer networks with a single neuron (Gunasekar et al. (2018)), as well as maximum-margin classification for logistic regression problems in separable data scenarios (Soudry et al. (2018); Chizat & Bach (2020)). Recent work by Li et al. (2022) identifies sufficient conditions under which the reparametrization $\phi$ makes the flow a mirror flow, requiring that the parametrization be commuting. However, this commuting condition is rarely satisfied in practical scenarios. A weaker condition, introduced by Marcotte et al. (2023), demands merely that the parametrization be involutive, thus ensuring that the flow on $z(t)$ can still be expressed as a Riemannian gradient flow (but not necessarly a mirror flow anymore) dependent solely on $z$ and the initial parameters $\theta(0)$. Marcotte et al. (2023) have shown that this weaker condition applies specifically to the parametrization used in two-layer ReLU networks (Stock & Gribonval (2022)).
Conservation laws
The functions conserved during the dynamics play a crucial role in rewriting flows in lower-dimensional representations. Indeed, when a trajectory $\theta(t)$ is constrained to remain within level sets $\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$ where $\mathbf{h}$ is a (vector-valued) conserved function, the dynamics are effectively restricted to a manifold of lower dimension. A particularly important class of conserved functions along these trajectories is given by the conservation laws associated with a certain architecture-dependent parametrization $\phi$, a concept introduced in Marcotte et al. (2023). These laws depend exclusively on $\phi$, and notably, in the context of neural network training dynamics, they represent quantities preserved across trajectories irrespective of the initial conditions or the training data-set. In the specific case of linear and ReLU neural networks, these conservation laws correspond exactly to previously known “canonical” conserved functions identified in Du et al. (2018), as demonstrated by Marcotte et al. (2023). Furthermore, Marcotte et al. (2023) establish that if the parametrization $\phi$ is involutive, there exist sufficiently many scalar conservation laws to fully rewrite the original trajectory $\theta(t)$ in terms of $\phi(\theta(t))$ and the initial conditions alone. In the linear network case, when so-called balanced conditions (a notion introduced in Arora et al. (2019)) are satisfied (i.e., when the initialization sets all canonical conservation laws (Chitour et al. (2018)) to zero), it becomes possible to rewrite the flow in terms of $z=\phi(\theta)$, where $\phi$ corresponds to the product of weight matrices, as a Riemannian metric that only depend on $z$ and $\mathbf{h}(\theta_{0})=0$ (Arora et al. (2018); Bah et al. (2022)). Moreover, Achour et al. (2025) extended this result to linear convolutional networks in the case of a mean squared loss, but this time for arbitrary initializations, with the Riemannian metric depending on the initialization. For linear networks and in the particular case when the loss function is the square loss, Bah et al. (2022) show that the trajectory evolves on the manifold of matrices having some fixed rank under balanced condition. Still in the square-loss setting, and in the case of two-layer linear networks, Tarmoun et al. (2021); Braun et al. (2022); Dominé et al. (2025) exploit the conservation laws to obtain an exact closed-form expression for $z(t)$ under specific configurations, whereas Varre et al. (2023) uses the same laws to analyse an implicit bias of this dynamic.
Our main contributions.
We first define the notion of intrinsic dynamic property (Definition˜2.6), then the notion of intrinsic metric property (Definition˜2.10) and finally the one of intrinsic recoverability property (Definition˜3.1), and we show the implications (Lemma˜2.11 and Lemma˜3.2):
$\text{Intrinsic Recoverability}\implies\text{Intrinsic Metric}\implies\text{Intrinsic Dynamic}.$
We then provide a simple criterion that characterizes the intrinsic recoverability property (Theorem˜3.3), and show (Proposition˜3.7) that this criterion is quasi equivalent to the Frobenius property (Definition˜3.6). We prove that the reparametrization for general ReLU networks of arbitrary depth satisfies this property (Theorem˜3.8), establishing that any initialization of a general ReLU network satisfies the intrinsic recoverability property (Corollary˜3.9), as illustrated by a characterization of the intrinsic dynamic of a 3-layer neural network (Proposition˜3.10). Next, by establishing a necessary condition for the intrinsic metric property to hold based on the study of kernels of linear mappings Theorem˜2.14, we show that the intrinsic metric property fails to hold for the natural reparametrizations corresponding of 2-layer linear networks (resp. of attention layers), unless the initialization satisfies the relaxed balanced condition introduced in Definition˜4.1 (Theorem˜4.4). We then show that relaxed balanced initializations do satisfy the intrinsic metric property, not only in 2-layer networks (Theorem˜4.3) but also in linear networks of arbitrary depth (Theorem˜4.6), and we characterize the resulting intrinsic dynamic. Finally, we extend our analysis to the infinite-depth limit of linear networks. We show that a set of functions is conserved along the trajectory (Proposition˜4.7), and, in contrast to the case $L>2$-layer, we derive a closed-form expression for the metric in the case of relaxed balanced initializations (Theorem˜4.8).
2 Dynamics of over-parameterized models
In most machine learning models, overparameterization occurs due to inherent symmetries (such as rescaling) within the parameter space $\theta\in\mathbb{R}^{D}$. In practice, this redundancy can be factored out through a function $\phi$ (often called a lifting (Candès et al., 2013; Gonon et al., 2023)) that captures these symmetries. Although the resulting lifted variable $z=\phi(\theta)\in\mathbb{R}^{d}$ often lives in higher dimension $d\gg D$, it also belongs to a lower dimensional manifold $\mathcal{Z}$ of dimension $d^{\prime}<D$, and provides a representation of the essential structure of the model.
We consider parameters $\theta(t)\in\mathbb{R}^{D}$ that evolve according to the gradient flow dynamic:
$\dot{\theta}(t)=-\nabla\ell(\theta(t)),$ (1)
to minimize the function $\ell$. In machine learning, $\ell(\theta)$ is typically defined as the empirical average over training samples $(x_{i},y_{i})$ of a quantity that depends on the output $g(\theta,x_{i})$ of a neural network with weights and biases collected in the parameter vector $\theta$. The function $g(\theta,x)$ can often be locally reparameterized via an architecture-dependent lifting $\phi(\theta)$, leading to the same factorization for the global loss $\ell$. This is the starting point of our analysis, captured via the following assumption:
Assumption 2.1 (Local reparameterization).
There exists $d$ and $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ such that: for each parameter $\theta_{0}\in\mathbb{R}^{D}$, there is an open neighborhood $\Omega\ni\theta_{0}$ and a function $f\in\mathcal{C}^{2}(\Omega,\mathbb{R})$ such that
$\forall\theta\in\Omega,\quad\ell(\theta)=f(\phi(\theta)).$ (2)
The following examples illustrate common choices of $\phi$ for various neural network architectures.
Example 2.2 (Linear neural networks).
For a two-layer network with $r$ hidden neurons and $\theta=(U,V)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m\times r}$ (where $D=(n+m)r$), the model $g(\theta,x)\coloneqq UV^{\top}x$ is factorized via the map $\phi_{\mathtt{Lin}}(\theta)\coloneqq UV^{\top}\in\mathbb{R}^{n\times m}$, thus the empirical risk $\ell$ can also be factorized by $\phi_{\mathtt{Lin}}$. This extends to $L$ layers where $\theta=(U_{L},\cdots,U_{1})$, with $g(\theta,x)\coloneqq U_{L}\cdots U_{1}x$ and $\phi_{\mathtt{Lin}}(\theta)\coloneqq U_{L}\cdots U_{1}$. The resulting factorization of $\ell$ holds globally on $\Omega=\mathbb{R}^{D}$.
Example 2.3 (ReLU neural networks).
Consider $g(\theta,x)=U\sigma(V^{\top}x)$, with $\sigma(y)\coloneqq(\max(y_{i},0))_{i}$ the ReLU activation function. Denoting $\theta=(U,V)$ with $U=(u_{1},\cdots,u_{r})\in\mathbb{R}^{n\times r}$, $V=(v_{1},\cdots,v_{r})\in\mathbb{R}^{m\times r}$ (so that $D=(n+m)r$). Consider $\theta(0)=(U(0),V(0))\in\mathbb{R}^{D}$ and let $x\in\mathcal{X}_{\theta(0)}:=\mathbb{R}^{m}-\cup_{j}\{{v_{j}(0)}^{\top}x=0\}$. Then on any domain $\Omega\subset\mathbb{R}^{D}$ such that $\theta(0)\in\Omega$ and $\theta\mapsto 1(v_{j}^{\top}x>0)=\epsilon_{j,x}$ is constant over $\theta\in\Omega$, the model $g_{\theta}(x)$ can be factorized by the reparametrization $\phi_{\mathtt{ReLU}}(\theta)=(u_{j}v_{j}^{\top})_{j=1}^{r}$ (here $d=rmn$) using $g(\theta,x)=\sum_{j}\epsilon_{j,x}\phi_{j}x$, so $\ell$ can be factorized by $\phi_{\mathtt{ReLU}}$ with some forward function $f$: the reparametrization $\phi_{\mathtt{ReLU}}(\theta)$ contains $r$ matrices of size $n\times m$ (each of rank at most one, so in particular one has $d^{\prime}\leq D-r$) associated to a “local” $f$ valid in a neighborhood of $\theta$. A similar factorization is possible for deeper ReLU networks (cf Neyshabur et al. (2015); Stock & Gribonval (2022); Gonon et al. (2023)) and we still write it $\phi_{\mathtt{ReLU}}$, as further discussed in the proof of Theorem˜3.8.
Example 2.4 (Attention layer).
For an attention layer, the input $X\in\mathbb{R}^{N\times\text{dim}}$ is the horizontal concatenation of $N$ tokens $x^{(i)}\in\mathbb{R}^{\textrm{dim}}$. The layer output is
$g(\theta,X)=\mathrm{softmax}(XQ^{\top}KX^{\top})XV^{\top}O\in\mathbb{R}^{N\times\textrm{dim}}\quad\text{where}\quad\mathrm{softmax}(A)_{i}=\frac{\exp(A_{i})}{\sum_{k=1}^{N}\exp(A_{ik})},$
with $Q,K,V,O\in\mathbb{R}^{d_{1}\times\text{dim}}$. We consider the reparameterization $\phi_{\mathtt{Att}}(\theta)\coloneqq(\phi_{1},\phi_{2})$ where $\phi_{1}\coloneqq Q^{\top}K$ and $\phi_{2}\coloneqq V^{\top}O$, such that $g(\theta,X)=\mathrm{softmax}(X\phi_{1}X^{\top})X\phi_{2}$, as done in Marcotte et al. .
Thus, similarly to the linear case Example˜2.2, $L$ can be globally factorized by $\phi_{\mathtt{Att}}$ as $f$ exhibits no dependency on the specific parameter configuration $\theta_{0}$. This naturally extends to multiple attention layers by concatenating the corresponding factorizations.
2.1 Dynamics of lifted parameters: to be or not to be intrinsic?
This paper addresses a fundamental question underlying much of the efforts to characterize the implicit bias of gradient-based methods: under what conditions does the gradient flow dynamics (1) in parameter space $\theta$ lead to a dynamics on the lifted parameters $z(t)\coloneqq\phi(\theta(t))$ that can be expressed as an intrinsic gradient flow on $z$? This is notably key when attempting to establish that $z(t)$ follows a mirror flow (Gunasekar et al., 2017), which is a key step to characterize the implicit bias of gradient-based optimization. We specifically examine when $z(t)$ follows a flow with respect to a Riemannian metric which, by definition depends only on $z$ (and the initial parameter configuration $\theta_{0}$), thereby eliminating explicit dependence on the parameter trajectory $\theta(t)$.
A starting point of the analysis is that, under ˜2.1 and by the chain rule
$\dot{z}(t)=\partial\phi(\theta(t))\dot{\theta}(t)=-\partial\phi(\theta(t))\partial\phi(\theta(t))^{\top}\nabla f(z(t)).$ (3)
Thus our goal is to understand when the symmetric, positive semi-definite matrix
$M(\theta):=\partial\phi(\theta)\partial\phi(\theta)^{\top}$ (4)
(corresponding to the so-called path kernel in when $\Phi$ is the path-lifting associated to ReLU networks Gebhart et al. (2021)) can be solely expressed in terms of $z$ and $\theta_{0}$ during the trajectory, i.e. do we have a function $K=K_{\theta_{0}}$ such that $M(\theta(t))=K(z(t))$? When this is the case (3) becomes
$\dot{z}(t)=-K(z)\nabla f(z),$ (5)
an ordinary differential equation (ODE) which is a Riemannian flow for the metric $K^{-1}(z)$ (or a sub-Riemannian flow for the pseudo-inverse $K^{+}(z)$ when $K(z)$ is not invertible) Boumal (2023), hence associated to an intrinsic dynamic on the lifted parameters $z(t)$.
As illustrated with the following example, rewriting $M(\theta(t))$ as a function of $z(t)$ along the trajectory $\theta(x)$ is indeed possible for simple linear networks, with a function $K(\cdot)$ that depends on the initialization $\theta_{0}$.
Example 2.5 (A simple linear network).
Consider $g(\theta,x)=uvx$, with $\theta\coloneqq(u,v)\in\mathbb{R}_{\star}\times\mathbb{R}^{m}$, and $z=\phi(\theta)=uv\in\mathbb{R}^{m}$ (cf Example˜2.2). Then $M(\theta)=\partial\phi(\theta)\partial\phi(\theta)^{\top}=vv^{\top}+u^{2}I_{m}$. During the trajectory $u^{2}-\|v\|^{2}=u_{0}^{2}-\|v_{0}\|^{2}=:\lambda$ (as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h(\theta):=}u^{2}-\|v\|^{2}$ is conserved (Arora et al. (2019)), and as $vv^{\top}=u^{-2}zz^{\top}$ we have $(u^{2})^{2}-\lambda u^{2}-\|z\|^{2}=0$ so that $u^{2}=\frac{\lambda+\sqrt{\lambda^{2}+4\|z\|^{2}}}{2}$. As a result along the whole trajectory we have $M(\theta)=K_{\theta_{0}}(z)$ so that $z(t)$ satisfies the ODE (5) with
$K_{\theta_{0}}(z)=\frac{2}{\lambda+\sqrt{\lambda^{2}+4\|z\|^{2}}}zz^{\top}+\frac{\lambda+\sqrt{\lambda^{2}+4\|z\|^{2}}}{2}\mathrm{Id}_{m},\quad\forall z.$
In particular when $m=1$ one has $K_{\theta_{0}}(z)=\sqrt{(u_{0}^{2}-v_{0}^{2})^{2}+4z^{2}}$ hence $\dot{z}=-\sqrt{\lambda^{2}+4z^{2}}\nabla f(z).$
See Appendix˜A for more comments on that example. In the above example, the function $K_{\theta_{0}}$, as its notation suggests, only depends on the initialization but not on the function $f$ such that $\ell=f\circ\phi$. In machine learning scenarios, $f$ typically captures dependence on the training dataset. The intrinsic metric $K_{\theta_{0}}(z)$ thus captures parts of the dynamics of $z(t)$ due to the network architecture (via $\phi$) and of the training algorithm (the gradient flow (1)) irrespective of the dataset and the learning task (of course the latter still play a role via the $\nabla f(z)$ term in the ODE $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$). This motivates the introduction of the following definition.
Definition 2.6 (Intrinsic dynamic property).
An initialisation $\theta_{0}$ verifies the intrinsic dynamic property on $\Omega$ with respect to $\phi$, if there is $K_{\theta_{0}}:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ such that, for any $f\in\mathcal{C}^{2}$, the maximal solution $\theta(\cdot)$ of (1) with $\ell=f\circ\phi$ satisfies $M(\theta(t))=K_{\theta_{0}}(\phi(\theta(t)))$ for each $t$ such that $\theta(t)\in\Omega$.
2.2 Conservation laws
Example˜2.5 illustrates a phenomenon that we will systematically exploit in our analysis: with the typical reparameterizations $\phi$ mentioned above, there exists a vector-valued function $\mathbf{h}:\theta\mapsto\mathbf{h}(\theta)\in\mathbb{R}^{N}$ that is conserved along the trajectory and allows to exhibit a function $K_{\theta_{0}}$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ along the trajectory. As these will play a key role in our analysis we now introduce the essential concepts related to conservation laws.
We denote $\phi_{1},\cdots,\phi_{d}$ the $d$ coordinate functions of the reparameterization $\phi:\mathbb{R}^{D}\mapsto\phi(\theta)\in\mathbb{R}^{d}\in\mathcal{C}^{\infty}(\mathbb{R}^{D},\mathbb{R}^{d})$. Since $\phi$ yields a factorization of the loss, functions $h$ such that $\nabla h(\theta)\perp\nabla\phi_{i}(\theta)$ for each $i$ and each $\theta$ remain constant along the trajectory. This has been thoroughly analyzed, see e.g. Marcotte et al. (2023, 2024), using the following definition.
Definition 2.7 (Conservation law for $\phi$).
A function $h\in\mathcal{C}^{1}(\Omega,\mathbb{R})$ is a conservation law for $\phi$ on $\Omega$ if for any $\theta\in\Omega$ one has $\partial\phi(\theta)\nabla h(\theta)=0$, i.e. for each $\theta\in\Omega$ and $i$, $\langle\nabla\phi_{i}(\theta),\nabla h(\theta)\rangle=0$.
Proposition 2.8.
If $\phi$ satisfies ˜2.1. and $h\in\mathcal{C}^{1}(\Omega,\mathbb{R})$ satisfies $\partial\phi(\theta)\nabla h(\theta)=0$ for every $\theta\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega}$, then $h$ remains constant during the trajectory $\theta(t)$ of (1) for any initialization $\theta_{0}\in\Omega$.
The conservation laws associated with a given parameterization $\phi$ have been almost exhaustively studied for parameterizations corresponding to linear networks, ReLU networks, and attention layers. In particular, prior work has shown that all conservation laws for $\phi$ in the cases of ReLU (cf Example˜2.3) and linear (cf Example˜2.2) networks (see Marcotte et al. (2023)) as well as for an attention layer (see Marcotte et al. ) are captured by the following proposition (Marcotte et al. (2023)). This has been proven theoretically for two-layer networks and empirically validated for deeper architectures using symbolic computation (see Marcotte et al. (2023)). It is worth noticing that all conservation laws in such cases are polynomials.
Proposition 2.9 (Conservation laws for classical $\phi$ on $\mathbb{R}^{D}$).
Consider $\theta=(U_{L},\cdots,U_{1})$ and $\phi_{\mathtt{Lin}}(\theta)\coloneqq U_{L}\cdots U_{1}$ from Example˜2.2 (resp. $\phi_{\mathtt{ReLU}}$ from Example˜2.3). The functions
$\mathbf{h}_{i}:\theta\mapsto U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}\,\text{(resp. }\mathbf{h}_{i}:\theta\mapsto\mathrm{Diag}(U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top})\text{)}$
are conservation laws for $\phi_{\mathtt{Lin}}$ (resp. $\phi_{\mathtt{ReLU}}$). Similary, considering $\theta\coloneqq(Q,K,V,O)$ and $\phi_{\mathtt{Att}}$ from Example˜2.4, $\mathbf{h}:\theta\mapsto(QQ^{\top}-KK^{\top},VV^{\top}-OO^{\top})$ is a set of conservation laws for $\phi_{\mathtt{Att}}$.
2.2.1 Intrinsic dynamics via conservation laws
Given conservation laws $\mathbf{h}(\theta)$ for $\phi$, the trajectory $\theta(t)$ for (1) remains at all times on the set
$\mathcal{M}_{\theta_{0}}\coloneqq\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\},$ (6)
determined by the initialization $\theta_{0}$. This holds true for any function $f$ such that $\ell=f\circ\phi$ (hence, in machine learning: for any task/loss and any dataset, provided that the neural network model is (locally) factorized via $\phi$).
To establish the existence of a function $K_{\theta}(\cdot)$ such that $M(\theta(t))=K_{\theta_{0}}(z(t))$ on the whole trajectory, a natural relaxation is thus to establish a related equality on the whole set $\mathcal{M}_{\theta_{0}}$ rather than only on a specific trajectory. This leads to the following definition and its immediate consequence.
Definition 2.10 (Intrinsic metric property).
We say that $\theta_{0}$ verifies the intrinsic metric property on an open set $U\ni\theta_{0}$ with respect to $\phi$, if there exists conservation laws $\mathbf{h}(\theta)\in\mathbb{R}^{N}$ for $\phi$ and a function $K_{\theta_{0}}\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}^{d\times d})$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}\cap U$.
Lemma 2.11.
If $\theta_{0}$ verifies the intrinsic metric property 2.10 on $U$ with respect to $\phi$, then it also verifies the intrinsic dynamic property 2.6 on $U$ with respect to $\phi$.
Remark 2.12.
It is not difficult to check on all examples considered in this paper that if $\theta_{0}$ satisfies the intrinsic metric property with respect to $\phi$ on some open set $U$, then any $\theta^{\prime}_{0}\in\mathcal{M}_{\theta_{0}}$ also satisfies the property on a properly modified open set $U^{\prime}$, with the same function $K$. This function thus only depends on $\mathbf{h}(\theta_{0})$, and we denote it $K_{\mathbf{h}(\theta_{0})}$ when needed to highlight this fact.
Remark 2.13.
Lemma˜2.11 remains valid with a slightly weakened version of Definition˜2.10, where $K_{\theta_{0}}$ is not required to be smooth. Yet, since the existence of a smooth solution to the resulting intrinsic ODE (5) is simplified when $K_{\theta_{0}}$ is $\mathcal{C}^{1}$ we chose to include this in the definition.
The following theorem (proved in Appendix˜B) establishes a necessary condition for the intrinsic metric property to hold. We use it to show that the property does not always hold for linear networks.
Theorem 2.14.
Consider $\mathbf{h}\in\mathcal{C}^{1}(\mathbb{R}^{D},\mathbb{R}^{N})$, $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$, and $\theta_{0}\in\mathbb{R}^{D}$ such that the matrix $\partial\mathbf{h}(\theta)\in\mathbb{R}^{N\times D}$ has constant rank on $\mathcal{M}_{\theta_{0}}\cap U$ with $U\ni\theta_{0}$ an open subset of $\mathbb{R}^{D}$ and $\mathcal{M}_{\theta_{0}}\coloneqq\mathbf{h}^{-1}(\{\mathbf{h}(\theta_{0})\})$. Then (i) $\implies$ (ii), where
  • (i)
    There exists an open set $O\supset\phi(\mathcal{M}_{\theta_{0}})\cap U$ and a map $K_{\theta_{0}}\in\mathcal{C}^{1}(O,\mathbb{R}^{d\times d})$ such that for each $\theta\in\mathcal{M}_{\theta_{0}}\cap U$: $M(\theta)=K_{\theta_{0}}(\phi(\theta))$;
  • (ii)
    For every $\theta\in\mathcal{M}_{\theta_{0}}\cap U$ one has
    $\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)\subseteq\mathrm{ker}\partial M(\theta).$ (7)
On one hand, a trivial case where (7) holds is when the intersection of kernels on the left hand side is zero:
$\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)=\{0\}.$ (8)
This stronger assumption can in fact be shown to imply the intrinsic metric property (see Theorem˜3.3 in the upcoming section), and we will show (cf Corollary˜3.9) that, with $\phi_{\mathtt{ReLU}}$ associated to general ReLU networks of any depth, there exists a set of conservation laws such that (8) indeed holds for any initialization. This implies the intrinsic metric property and therefore the intrinsic dynamic property irrespective of the initialization for ReLU networks with $\phi_{\mathtt{ReLU}}$.
On the other hand, for linear networks with more than one hidden neuron, we will show that it is not possible to reduce the problem to (8). Nevertheless, certain initializations (known as balanced conditions (Arora et al. (2019)) are known to satisfy the intrinsic metric property with respect to the reparametrisation $\phi_{\mathtt{Lin}}$ (cf (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2)). In this paper, we generalize this result to so-called relaxed balanced initializations (see Definition˜4.1). Moreover, we show that in certain configurations, relaxed balanced initializations are exactly the only ones that satisfy the intrinsic metric property (cf Theorem˜4.3 and Theorem˜4.4).
3 Intrinsic recoverability and application for ReLU neural networks
In this section we consider a stronger condition called intrinsic recoverability property which requires not only that $M(\theta(t))$ can be rewritten as a function of $z(t)$ and the initialization, but that at each point of the trajectory $\theta(t)$ itself can be fully expressed in terms of $z(t)$ and the initialization $\theta_{0}$. In other words, in this scenario, $\theta(t)$ can be completely recovered from the parameterization $\phi$ and the initialization alone, hence the name. As we will establish, this apparently strong property indeed holds when (8) is satisfied, which is always the case for ReLU networks.
3.1 Intrinsic recoverability implies intrinsic metric
Definition 3.1 (Intrinsic recoverability property).
We say that $\theta_{0}$ verifies the intrinsic recoverability property on an open set $U\ni\theta_{0}$ with respect to $\phi$, if there exists conservation laws $\mathbf{h}(\theta)\in\mathbb{R}^{N}$ for $\phi$ and a function $\Gamma(\cdot)\in\mathcal{C}^{1}(\mathbb{R}^{d}\times\mathbb{R}^{N},\mathbb{R}^{D})$ such that $\theta=\Gamma(\phi(\theta),\mathbf{h}(\theta))$ for each $\theta\in U$.
When this property holds, each $\theta\in\mathcal{M}_{\theta_{0}}$ satisfies $M(\theta)=M[\Gamma(\phi(\theta),\mathbf{h}(\theta))]=M[\Gamma(\phi(\theta),\mathbf{h}(\theta_{0}))]=K_{\mathbf{h}(\theta_{0})}(\phi(\theta))$ (with $K_{\mathbf{h}(\theta_{0})}(\cdot):=M[\Gamma(\cdot,\mathbf{h}(\theta_{0}))]$), hence the following result.
Lemma 3.2.
If $\theta_{0}$ satisfies the intrinsic recoverability property on an open set $U\ni\theta_{0}$ with respect to $\phi$, then $\theta_{0}$ satisfies the intrinsic metric property on $U$ with respect to $\phi$.
The intrinsic recoverability property is equivalent to (8) (see Appendix˜C for a proof):
Theorem 3.3.
Consider $\phi\in\mathcal{C}^{2}(\mathbb{R}^{D},\mathbb{R}^{d})$ and $\theta_{0}\in\mathbb{R}^{D}$. The following are equivalent: (i) there exists conservation laws $\mathbf{h}\in\mathcal{C}^{1}(\Omega,\mathbb{R}^{N})$ for $\phi$ on a neighborhood $\Omega$ of $\theta_{0}$ such that (8) holds for each $\theta\in\mathcal{M}_{\theta_{0}}\cap\Omega$; (ii) there exists an open set $U\subseteq\Omega$ on which $\theta_{0}$ satisfies the intrinsic recoverability property Definition˜3.1 (and thus the intrinsic metric property Definition˜2.10) with respect to $\phi$.
3.2 The Frobenius property is almost equivalent to intrinsic recoverability
We are interested in establishing condition (8), as it implies the intrinsic recoverability property, and thus an intrinsic dynamics. It may not seem obvious a priori how to verify whether such a condition can hold, nor how to construct suitable conservation laws $\mathbf{h}$ in practice. Intuitively, one should select as many conservation laws as possible while ensuring they remain independent, in a specific sense defined by (Marcotte et al., 2023, Definition 2.18). As shown by Marcotte et al. (2023), knowing the maximal number of such conservation laws can be checked using Lie brackets of the associated vector fields. We recall the relevant definitions and explain how this criterion applies in our setting.
Definition 3.4 (Lie brackets).
Given two vector fields $\chi_{1},\chi_{2}\in\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the Lie brackets $\chi_{1},\chi_{2}]$ is the vector field defined by $\chi_{1},\chi_{2}](\theta)\coloneqq\partial\chi_{2}(\theta)\chi_{1}(\theta)-\partial\chi_{1}(\theta)\chi_{2}(\theta)$.
Definition 3.5 (Generated Lie algebra).
Given some function space $\mathbb{W}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$, the generated Lie algebra of $\mathbb{W}$ is the smallest subspace of $\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$ that contains $\mathbb{W}$ and that is stable by Lie brackets, and is denoted $\mathrm{Lie}(\mathbb{W})$.
The trace at $\theta\in\Theta$ of any set $\mathbb{W}\subset\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{D})$ of vector fields is defined as the linear space
$\mathbb{W}(\theta)\coloneqq\mathrm{span}{\{\chi(\theta):\chi\in\mathbb{W}\}}\subseteq\mathbb{R}^{D},$ (9)
and for any infinitly smooth $\phi$ we denote $\mathbb{W}_{\phi}\coloneqq\mathrm{span}\{\nabla\phi_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\cdot),\ 1\leq i\leq d}\}\subseteq\mathcal{C}^{\infty}(\Theta,\mathbb{R}^{d})$.
Definition 3.6 (Frobenius property).
An infinitely smooth function $\phi$ satisfies the Frobenius property on $\Omega$ if for all $\theta\in\Omega$, $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)=\mathbb{W}_{\phi}(\theta)$.
The following proposition (see Appendix˜D for a proof) shows how this property implies the intrinsic dynamic property of $\theta_{0}$. In particular, as Frobenius property does not hold for $\phi_{\mathtt{Lin}}$ (Marcotte et al., 2023, Proposition I.1), it is not possible to have the intrinsic recoverability for linear networks with classical $\phi_{\mathtt{Lin}}$.
Proposition 3.7.
We have the following implications $\implies(ii)\implies(iii): (i)(i)$ $\phi$ satisfies the Frobenius property on $\Omega$ and the trace of $\mathbb{W}_{\phi}$ has its dimension that is constant on $\Omega$; $(ii)$ For any $\theta_{0}\in\Omega$, there exists conservation laws $\mathbf{h}$ for $\phi$ on a neighborhood $U\subset\Omega$ of $\theta_{0}$ such that for each $\theta\in\mathcal{M}_{\theta_{0}}$ (8) holds; $(iii)$ $\phi$ satisfies the Frobenius property on $\Omega$.
The result presented in Li et al. (2022) can be recovered as a special case of Proposition˜3.7: the authors require the reparametrization $\phi$ to be commuting, meaning that for all pairs $\phi_{i},\phi_{j}$, the Lie bracket $\nabla\phi_{i},\nabla\phi_{j}]$ is equal to zero. In this setting, $\phi$ naturally satisfies the Frobenius property, and the result of Li et al. (2022) establishes an even stronger property: the dynamics on $\phi(\theta)$ form a mirror flow. In particular, it is worth noting that diagonal networks satisfy that their parametrization (product of the diagonals) is commuting (as all coordinates functions are separable), which thus (Li et al. (2022)) implies a mirror flow dynamic, and thus an implicit bias (see e.g. Azulay et al. (2021)). In contrast, we consider a weaker condition; we seek only to determine whether the dynamics on $z=\phi(\theta)$ can be expressed intrinsically as a Riemannian gradient flow.
3.3 Application to general deep ReLU networks
We now show that the intrinsic recoverability property is satisfied for any initialisation for the parametrisation $\phi$ associated to a large class of (deep) ReLU neural networks. While this result is already known in the two-layer case (Marcotte et al., 2023, Examples 3.5 and 3.8), here we establish it for the general model of ReLU networks of Gonon et al. (2023), associated to a directed acyclic graph (DAG) of any depth, including skip connections and arbitrary mixes of ReLU/linear/max-pooling activations. We first establish that $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property (see Appendix˜E for a proof):
Theorem 3.8.
[…]
This leads to the following corollary (proved in Appendix˜F) which guarantees the existence of a maximal set of conservation laws big enough to ensure the intrinsic recoverability property.
Corollary 3.9.
There exists a dense open set $\Theta$ of $\mathbb{R}^{D}$ such that any $\theta_{0}\in\Theta$ admits an open neighborhood $U\subseteq\Theta$ on which $\theta_{0}$ satisfies the intrinsic recoverability property, and thus the intrinsic dynamic property with respect to $\phi_{\mathtt{ReLU}}$.
In practice, the known conservation laws given in Proposition˜2.9 yield, on a dense open subset, $m$ independent conservation laws, where $m$ corresponds to the number of hidden neurons. To verify that these are indeed the only ones, one must check that the trace of $\mathbb{W}_{\phi_{\mathtt{ReLU}}}$ has dimension DmD-m; while we do not prove this here, it is empirically supported by Marcotte et al. (2023), which confirms that $\mathrm{Lie}\big{(}\mathbb{W}_{\phi_{\mathtt{ReLU}}}\big{)}(\theta)=\mathbb{W}_{\phi_{\mathtt{ReLU}}}(\theta)$ has dimension DmD-m when sampling random values of $\theta$, as well as random dimensions and depths.
As a concrete example the following proposition (proved in Appendix˜G) provides the first closed form formula of the intrinsic dynamic for a three-layer ReLU network with scalar input and output.
Proposition 3.10.
For a 3-layer ReLU MLP with scalar input/output, the factorization $\phi_{\mathtt{ReLU}}$ reads [^1] When written as a $n\times m$ matrix, we denote $Z$ instead of $z$ and also view $\nabla f(Z)$ as an $n\times m$ matrix.
$Z=\phi_{\mathtt{ReLU}}(u,V,w):=\mathrm{diag}(u)\,V\,\mathrm{diag}(w)\in\mathbb{R}^{n\times m},$
with $u\in\mathbb{R}^{n}$, $V\in\mathbb{R}^{n\times m}$, and $w\in\mathbb{R}^{m}$. Define $\Theta:=\{(u,V,w):u_{i},V_{ij},w_{j}\neq 0\ \forall i,j\}$, and let $\theta(t)$ be the maximal solution to (1) with $\theta(0)=\theta_{0}\in\Theta$. The flow preserves the $n+m$ conservation laws $\mathbf{h}(\theta):=\big{(}(u_{i}^{2}-\sum_{j}V_{ij}^{2})_{i=1}^{n},(w_{j}^{2}-\sum_{i}V_{ij}^{2})_{j=1}^{m}\big{)}$, and the intrinsic dynamics $\dot{z}=-K_{\theta_{0}}(z)\nabla f(z)$ on $z=\mathtt{vec}(Z)$ corresponds to
$\dot{Z}=-\mathrm{ddiag}(\nabla f(Z)Z^{\top})\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\,Z-\mathrm{diag}(\boldsymbol{\alpha})\,\nabla f(Z)\,\mathrm{diag}(\boldsymbol{\beta})-Z\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\,\mathrm{ddiag}(Z^{\top}\nabla f(Z)),$
where: a) for any matrix $M$, we denote $\mathrm{ddiag}(M)=\mathrm{diag}\bigl{(}\mathrm{Diag}(M)\bigr{)}$, where $\mathrm{Diag}(M)$ extracts its diagonal as a vector and $\mathrm{diag}(v)$ denotes the diagonal matrix with entries of $v$; and b) the vectors $\boldsymbol{\alpha}=\boldsymbol{\alpha}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{n}$ and $\boldsymbol{\beta}:=\boldsymbol{\beta}(Z,\mathbf{h}(\theta_{0}))\in\mathbb{R}_{>0}^{m}$ (uniquely determined by $Z$ and $\mathbf{h}(\theta_{0})$) satisfy
$\boldsymbol{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{n}-\boldsymbol{\lambda}\odot\boldsymbol{\alpha}=0,\quad\boldsymbol{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{m}-\boldsymbol{\mu}\odot\boldsymbol{\beta}=0,$ (10)
with $|Z|^{2}\in\mathbb{R}^{n\times m}$ the element-wise square on the matrix $Z\in\mathbb{R}^{n\times m}$) and with $\boldsymbol{\lambda}\in\mathbb{R}^{n}$, $\boldsymbol{\mu}\in\mathbb{R}^{m}$ such that $\mathbf{h}(\theta_{0})=(\boldsymbol{\lambda},\boldsymbol{\mu})$. When $\boldsymbol{\lambda},\boldsymbol{\mu}=0$, (10) entirely characterizes $(\boldsymbol{\alpha},\boldsymbol{\beta})$.
4 Deep linear neural networks and linear neural ODEs
For $L$-layer linear networks, $\theta=(U_{1},\ldots,U_{L})$ and the path-lifting formalism (Gonon et al., 2023) yields a factorization via $\phi_{\mathtt{ReLU}}$, leading to an intrinsic dynamics by the results of the previous section. It is more common however to consider the dynamics of $Z_{L}:=\phi_{\mathrm{Lin}}(\theta_{L})=U_{L}\cdots U_{1},$ since $\phi_{\mathtt{lin}}$ is more efficient that $\phi_{\mathtt{ReLU}}$ in terms of dimension reduction. We now analyze the dynamics of $Z_{L}(t)$. The gradient flow $\dot{\theta}_{L}=-\nabla\ell(\theta_{L})$ gives the evolution of $Z_{L}$ (see e.g. (Bah et al., 2022, Lemma 2)):
$\dot{Z}_{L}=-\sum_{j=1}^{L}S_{j}\,\nabla f(Z_{L})\,T_{j},\quad\text{with}\quad\left\{\begin{aligned} S_{j}&:=U_{L}\cdots U_{j+1}\,U_{j+1}^{\top}\cdots U_{L}^{\top},\quad&&S_{L}=\mathrm{Id},\\ T_{j}&:=U_{1}^{\top}\cdots U_{j-1}^{\top}\,U_{j-1}\cdots U_{1},\quad&&T_{1}=\mathrm{Id}.\end{aligned}\right$. (11)
Observe that the metric $M(\theta_{L})$ on $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z_{L}=\mathtt{vec}(Z_{L})$} is entirely characterized by the functions
$(S_{j}(\theta_{L}),T_{j+1}(\theta_{L}))_{j=1}^{L-1}.$ (12)
Definition 4.1 (Relaxed balanced conditions).
We say that $\theta_{L}\coloneqq(U_{L},\cdots,U_{1})$ satisfies the relaxed balanced condition if there exists $\boldsymbol{\lambda}\coloneqq(\lambda_{i})_{i}\in\mathbb{R}^{L-1}$ such that
$U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}={\lambda_{i}}\mathrm{Id},\quad\forall 1\leq i\leq L-1.$ (13)
(0-)balanced conditions (Bah et al., 2022, Def 1) (Arora et al., 2019, Def 1) correspond to $\boldsymbol{\lambda}=0$.
Remark 4.2.
It is worth noting that Dominé et al. (2025) used this exact same condition and called it the $\lambda$-balanced condition. However, the definition of $\lambda$-balanced condition is already used (see (Arora et al., 2019, Def 1)) by the literature to refer to the weaker condition $\|U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}\|\leq{\lambda_{i}}.$ Other works (see e.g. Tarmoun et al. (2021); Braun et al. (2022); Varre et al. (2023)) use stronger conditions on the initializations, that satisfy in particular the relaxed balanced conditions of Definition˜4.1.
4.1 Deep linear neural networks
We first detail the study of the two-layer case, and then generalize it to the deep case.
Matrix factorization. We consider the two-layer case where $\theta\coloneqq(U,V)\in\mathbb{R}^{n\times r}\times\mathbb{R}^{m\times r}$ and with $Z=\phi_{\mathtt{Lin}}(\theta)\coloneqq UV^{\top}\in\mathbb{R}^{n\times m}$. We assume $\theta(t)$ satisfies the gradient flow (1) with $\theta(0)=(U_{t=0},V_{t=0})$. We denote $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\in\mathbb{R}^{r\times r}$.
If $\theta_{0}$ satisfies the balanced condition (13) $S=0$, then (Arora et al., 2018, Theorem 1) (Bah et al., 2022, Lemma 2) $\theta_{0}$ satisfies the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$ and
$\dot{Z}=-\sqrt{ZZ^{\top}}\nabla f(Z)-\nabla f(Z)\sqrt{Z^{\top}Z}.$ (14)
We now generalize this result (see Appendix˜H for a proof) to a broader class of initializations: all initializations satisfying the relaxed balanced condition (13) possess the intrinsic metric property.
Theorem 4.3.
Consider $\theta_{0}\coloneqq(U_{t=0},V_{t=0})$ where both $U_{t=0}\in\mathbb{R}^{n\times r}$ and $V_{t=0}\in\mathbb{R}^{m\times r}$ have full rank $r\leq\text{min}(n,m)$, and assume $S=\lambda\mathrm{Id}_{r}$ for some $\lambda\in\mathbb{R}$. Then, on a neighborhood $\Omega$ of $\theta_{t=0}$:
$\dot{Z}=-\Pi_{ZZ^{\top}}\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(X)-\nabla f(X)\Pi_{Z^{\top}Z}\left[-\frac{\lambda}{2}\mathrm{Id}_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$ (15)
where $\Pi_{A}$ is the orthogonal projector on $\mathrm{range}{A}$.
Note that (14) corresponds indeed to (15) with $\lambda=0$. Note also that Theorem˜4.3 generalizes to the case $r\leq\text{min}(n,m)$ the expression obtained in (Dominé et al., 2025, Theorem 5.2) for the special case $r=\text{min}(n,m)$ (if Dominé et al. (2025) focus in general on the squared loss, the proof of their Theorem 5.2 does not rely on the use of this specific loss: this result can be applied for any loss, as ours).
The following theorem shows that the relaxed balanced condition is actually a necessary condition when $r\leq\text{max}(n,m)$ to have the intrinsic metric property. Its proof (see Appendix˜I) relies on showing the non-inclusion of the kernels of (7).
Theorem 4.4.
Let $\theta_{0}\coloneqq(U_{t=0},V_{t=0})$. Assume that both $U_{t=0}\in\mathbb{R}^{n\times r}$ and $V_{t=0}\in\mathbb{R}^{m\times r}$ have a full rank and that $r\leq\text{max}(n,m)$. If $S\coloneqq U_{t=0}^{\top}U_{t=0}-V_{t=0}^{\top}V_{t=0}\neq\lambda\mathrm{Id}_{r}$, then $\theta_{0}$ does not satisfy the intrinsic metric property (Definition˜2.10) with respect to $\phi_{\mathtt{Lin}}$.
The case when $r>\text{max}(n,m)$ is still an open question. In the particular case $n=m=1$, for any $r$, the following proposition (see Appendix˜J for a proof) shows that all initializations do satisfy the intrinsic metric property with respect to $\phi_{\mathtt{Lin}}$.
Proposition 4.5.
Let $\theta\coloneqq(u,v)$ with $u\in\mathbb{R}^{r}$ and $v\in\mathbb{R}^{r}$. Then $z\coloneqq\phi_{\mathtt{Lin}}(\theta)=\langle u,v\rangle\in\mathbb{R}$. We denote $S\coloneqq u_{t=0}u_{t=0}^{\top}-v_{t=0}v_{t=0}^{\top}\in\mathbb{R}^{r\times r}$. Then one has $\dot{z}=-\sqrt{2\mathrm{tr}(S^{2})-\mathrm{tr}(S)^{2}+4z^{2}}\nabla f(z).$
In particular, it is important to note that the two-layer linear analysis allows these results to be applied directly to networks composed of attention layers (Example˜2.4).
Deep linear neural networks. We now consider linear neural networks of arbitrary depth, with square weight matrices $\theta_{L}\coloneqq(U_{L},\dots,U_{1})$ where $U_{i}\in\mathbb{R}^{n\times n}$. The following theorem (see Appendix˜K for a proof) generalizes Theorem˜4.3 to this setting.
Theorem 4.6.
If $\theta_{L}(0)$ satisfies the relaxed balanced condition (Definition˜4.1) with $\boldsymbol{\lambda}=(\lambda_{i})_{i}$ then during the trajectory $\theta_{L}(t)$ of (1), the matrices in (11) satisfy $S_{j}(\theta_{L}(t))=Q_{j}(U_{L}(t)U_{L}(t)^{\top})$ and $T_{j}(\theta_{L}(t))=R_{j}(U_{1}(t)^{\top}U_{1}(t))$, where $Q_{j}(x):=\prod_{k=0}^{L-j-1}(x-a_{k})$ with $a_{0}:=0$ and $a_{k}:=\sum_{i=1}^{k}\lambda_{L-i}$ for $k=1,\cdots L-1$ and $R_{j}(x):=\prod_{k=0}^{j-2}(x-b_{k})$ with $b_{0}:=0$ and $b_{k}:=-\sum_{i=1}^{k}\lambda_{i}$. Moreover $U_{L}U_{L}^{\top}$ (resp. $U_{1}^{\top}U_{1}$) is the unique root of $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ (resp. of $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$) with spectrum lower bounded by $\max_{0\leq k\leq L-1}a_{k}$ (resp. by $\max_{0\leq k\leq L-2}b_{k}$). This implies that all matrices in (12) are entirely characterized by $Z_{L}$ and the initialization, hence $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
In the particular case of balanced conditions ($\boldsymbol{\lambda}=0$), our theorem recovers the dynamics described in (Arora et al., 2018, Theorem 1), (Bah et al., 2022, Lemma 2).
Network type Parametrization $\boldsymbol{\phi_{\text{Lin}}}$ $\boldsymbol{\phi_{\text{ReLU}}}$
Linear network IMP: only for relaxed balanced $\theta_{0}$ Dimension: $d^{\prime}\boldsymbol{\ll}D$ IMP: for all $\theta_{0}$ $d^{\prime}=D-k\boldsymbol{\approx}D$
DAG ReLU N/A IMP: for all $\theta_{0}$ $d^{\prime}=D-k\boldsymbol{\approx}D$
IMP = intrinsic metric property ; $d^{\prime}$ = dimension of the manifold of $\phi(\theta)$; kk= # hidden neurons
Table 1: The table summarizes which parametrizations can be used to analyze which type of neural network, along with the corresponding results.
4.2 Infinitely deep linear networks
We next consider the limit when $L\rightarrow+\infty$ of deep linear residual networks with parameters $U_{k}=\mathrm{Id}_{n}+\mathcal{A}_{\frac{k}{L}}$, and thus focus on the analysis of the parameter $\theta=(\mathcal{A}_{s})_{s\in[0,1]}$, where $\mathcal{A}_{s}\in\mathbb{R}^{n\times n}$: this corresponds to studying linear neural ODEs (introduced by Chen et al. (2018)). Remarkably, our theoretical approach continues to apply in this regime, and yields a closed-form formula for the metric. In this section, we thus study the dynamics of parameters $\theta(t)\in\mathcal{X}$ where $\mathcal{X}$ corresponds to the Banach space $(\mathcal{C}^{1}([0,1],\mathbb{R}^{n\times n}),\|\cdot\|_{\mathcal{C}^{1}})$ where $\|f\|_{\mathcal{C}^{1}}:=\max\!\bigl{\{}\|f\|_{\infty},\,\|f^{\prime}\|_{\infty}\bigr{\}}$, and such that the trajectory $t\mapsto\theta(t)=(\mathcal{A}_{s}(t))_{s\in[0,1]}$ is the solution of the gradient flow on $\ell(\theta)$, given by the (family of coupled) ODE
$\forall s\in[0,1],\quad\frac{\partial\mathcal{A}_{s}}{\partial t}(t)=-\,\mathfrak{g}_{s}(t),\quad\text{with}\quad\mathfrak{g}_{s}(t):=\frac{\partial\ell}{\partial\mathcal{A}_{s}}\Bigl{(}\theta(t)\Bigr{)}\in\mathbb{R}^{n\times n},$ (16)
where we assume that the loss function $\ell:\mathcal{X}\mapsto\mathbb{R}$ is such that $\theta\mapsto(\frac{\partial\ell}{\partial\mathcal{A}_{s}}(\theta))_{s\in[0,1]}$ is locally Lipschitz on $\mathcal{X}$ (to ensure by the Cauchy-Lipschitz theorem that indeed there exists a unique maximal solution $\theta(\cdot)\in\mathcal{C}^{1}([0,T),\mathcal{X})$ of (16) with a given $\theta(0)$).
Now as an infinite-depth analog of the product $Z_{L}=U_{L}\ldots U_{1}$, given any $\theta\in\mathcal{X}$ we consider $s\in[0,1]\mapsto Z_{s}=Z_{s}[\theta]\in\mathbb{R}^{n\times n}$ the unique global solution (as $\theta=(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}$) of the state equation
$\frac{\mathrm{d}}{\mathrm{d}s}Z_{s}=\mathcal{A}_{s}Z_{s},\quad Z_{0}=\mathrm{Id}_{n}.$ (17)
The analog to ˜2.1 is to assume that $\ell(\theta)=f\bigl{(}Z_{s=1}\bigr{)}$ with $f\in\mathcal{C}^{1}$, and we now want to know if it is possible to rewrite the dynamic $\frac{\partial Z_{s=1}}{\partial t}(t)$ as an intrinsic dynamic that only depends on $Z_{s=1}(t)$ and the initialization $\theta(0)$.
The following proposition (see Appendix˜L for a proof) gives a set of conserved functions during all trajectories of (16).
Proposition 4.7.
For any $\theta\coloneqq(\mathcal{A}_{s})_{s\in[0,1]}\in\mathcal{X}$, we define for any $s\in[0,1]$ the function $\mathbf{h}_{s}:\theta\mapsto\mathcal{A}_{s}^{\prime}+\mathcal{A}_{s}^{\prime\top}+[\mathcal{A}_{s}^{\top},\mathcal{A}_{s}]\in\mathbb{R}^{n\times n},$ where we denote $\mathcal{A}_{s}^{\prime}\coloneqq\frac{\mathrm{d}}{\mathrm{d}s}\mathcal{A}_{s}$. Then for any $s\in[0,1]$, one has for any $t$: $\mathbf{h}_{s}(\theta(t))=\mathbf{h}_{s}(\theta(0))$, where $\theta(t)$ is the maximal solution of (16) with initialization $\theta(0)$.
Moreover, the following theorem (see Appendix˜M for a proof) shows that for relaxed balanced initializations, the evolution of $Z_{1}(t)=Z_{s=1}(t)$ is entirely described by $Z$ and the initialization.
Theorem 4.8.
If the initialization $\theta(0)$ satisfies that for each $s\in[0,1]$ $\mathbf{h}_{s}(\theta(0))=\lambda(s)\mathrm{Id}_{n}$ for some $\lambda(\cdot)\in\mathcal{C}^{0}([0,1],\mathbb{R})$, then one has
$\dot{Z}_{1}=-\int_{0}^{1}(Z_{1}Z_{1}^{\top})^{1-s}\exp(\gamma(s))\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s}\mathrm{d}s,$
with $\gamma(s)\coloneqq(1-s)\psi_{1}(1)-\psi_{1}(1-s)-s\psi_{2}(1)+\psi_{2}(s)$, where $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$ and $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. In particular, if $\lambda(\cdot)\equiv 0$ (balanced-condition), then $\gamma(\cdot)\equiv 0$.
In a sense, this theorem captures the infinite-depth limit ($L\rightarrow+\infty$) of Theorem˜4.6, while offering the key advantage of an explicit closed-form expression for the associated metric.
Conclusion
In this paper, we investigated when high-dimensional gradient flows can be recast as intrinsic Riemannian flows in lower-dimensional spaces. Our results show that such reductions are always possible for ReLU networks under path-lifting parametrization, and for linear networks under relaxed balanced initializations. A central contribution is our analysis of the “path-lifting metric”, a recently introduced and still largely unexplored object, for which we provide an intrinsic characterization in the 3-layer case. Extending this analysis to deeper or more general architectures could shed new light on the geometry of gradient dynamics for general ReLU networks.
Aknowledgments
The work of G. Peyré was supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir" program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and by the European Research Council (ERC project WOLF). The work of R. Gribonval was supported by the AllegroAssai ANR-19-CHIA-0009 project of the French Agence Nationale de la Recherche and by the SHARP ANR project ANR-23-PEIA-0008 in the context of the France 2030 program.
Appendix A More comments on Example˜2.5
In example (2.5), when $\lambda\rightarrow+\infty$ $K_{\theta_{0}}/\lambda\to I_{m}$ (the Euclidean metric). When $\lambda=0$, one has $K_{\theta_{0}}(z)=\|z\|I_{m}+\frac{zz^{\top}}{\|z\|}$ for every $z\neq 0$, and in particular, by the uniqueness result in the Cauchy-Lipschitz theorem, 0 is reachable only if $z_{0}=0$. When $\lambda\rightarrow$-$\infty$, $K_{\theta_{0}}/|\lambda|\to\frac{zz^{\top}}{\|z\|^{2}}$. See https://github.com/sibyllema/Intrinsic_training_dynamic for numerical illustrations of these different behaviors.
Appendix B Proof of Theorem˜2.14
See 2.14
Proof.
(i) $\Rightarrow$ (ii).
Assume (i) and fix $\theta\in\mathcal{M}_{\theta_{0}}\cap U$ and a vector $v\in\ker\partial\phi(\theta)\cap\ker\partial\mathbf{h}(\theta)$. Applying the chain rule in the ambient space $\mathbb{R}^{D}$ (possible as $v\in\ker\partial\mathbf{h}(\theta)=T_{\theta}\mathcal{M}_{\theta_{0}}$ because $\partial\mathbf{h}(\theta)$ has its rank that remains constant on $\mathcal{M}_{\theta_{0}}\cap U$ by hypothesis) gives
$\partial M(\theta)\cdot v=\partial K_{\theta_{0}}(\phi(\theta))\cdot\bigl{(}\partial\phi(\theta)\cdot v\bigr{)}=\partial K_{\theta_{0}}(\theta)\cdot 0=0,$
hence $v\in\ker\partial M(\theta)$ and (ii) holds. ∎
Appendix C Proof of Theorem˜3.3
See 3.3
Proof.
We first show $(i)\implies(ii)$. We assume $(i)$. Observe that given any $\theta\in\mathbb{R}^{D}$(8) is equivalent (by rank theorem) to
$\mathrm{rank}\begin{pmatrix}\partial\phi(\theta)\\ \partial\mathbf{h}(\theta)\end{pmatrix}=D.$ (18)
By smoothness of $\phi$ and $\mathbf{h}$, if (18) holds at $\theta_{0}$ is also holds in a whole neighborhood $U$ of $\theta_{0}$. By the implicit function theorem, denoting $F(\theta)\coloneqq(\phi(\theta),\mathbf{h}(\theta))$, it implies that $\theta$ can be expressed on $UU$ as $\theta=F^{-1}(\phi(\theta),\mathbf{h}(\theta))=\Gamma(\phi(\theta),\mathbf{h}(\theta))$.
We now show $(ii)\implies(i)$. We assume $(ii)$. Then on $U$ one has $\theta=\Gamma(\phi(\theta),\mathbf{h}(\theta))$. Thus on $\mathcal{M}_{\theta_{0}}\cap U$, one has $\theta=\Gamma(\phi(\theta),\mathbf{h}(\theta_{0}))$. We now fix some $\theta\in\mathcal{M}_{\theta_{0}}\cap U$ and we consider a vector $v\in\ker\partial\phi(\theta)\cap\ker\partial\mathbf{h}(\theta)$.
Applying the chain rule in the ambient space $\mathbb{R}^{D}$ on $\Gamma$ gives
$v=\mathrm{Id}_{D}v=\partial_{\phi(\theta)}\Gamma(\phi(\theta),\mathbf{h}(\theta_{0}))\cdot\bigl{(}\partial\phi(\theta)\cdot v\bigr{)}=\partial\Gamma(\theta)\cdot 0=0,$
and thus $v=0$. ∎
Appendix D Proof of Proposition˜3.7.
See 3.7
Proof.
$(i)\implies(ii)$ is direct consequence of the proof of (Marcotte et al., 2023, Proposition 3.7).
We now show $(ii)\implies(iii)$. Let us assume $(ii)$. We fix $\theta_{0}$. Then by assumption on $U\ni\theta_{0}$, $\theta=\Gamma(\mathbf{h}(\theta),\phi(\theta))$, and by using Proposition˜3.7 one has $\ker\partial\phi(\theta)\cap\ker\partial\mathbf{h}(\theta)=\{0\}$ on $\mathcal{M}_{\theta_{0}}\cap U$. Thus (18) holds on a open neighborhood $O$ of $\theta_{0}$. As $\mathbf{h}$ are conservation laws for $\phi$ on $U$, one has on $O\cap U$ that $\mathbb{W}_{\phi}(\theta)=D-\text{rank}\partial h(\theta)$.
But as $\langle\nabla h(\theta),\nabla\phi_{i}(\theta)\rangle=\langle\nabla h(\theta),\nabla\phi_{j}(\theta)\rangle=0\implies\langle\nabla h(\theta),[\nabla\phi_{i},\nabla\phi_{j}](\theta)\rangle=0$, one has necessarly $\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})\leq D-\text{rank}\partial h(\theta_{0})$ and as one also has $D-\text{rank}\partial h(\theta_{0})=\mathrm{dim}\mathbb{W}_{\phi}(\theta_{0})\leq\mathrm{dim}\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$ then one has $\mathbb{W}_{\phi}(\theta_{0})=\mathrm{Lie}\mathbb{W}_{\phi}(\theta_{0})$. This holds for any $\theta_{0}$, which concludes the proof. ∎
Appendix E Proof of Theorem˜3.8.
See 3.8
Proof.
[…]
$\theta=\bigl{(}\theta_{i\cap j},\;\theta_{i\setminus j},\;\theta_{j\setminus i},\;\theta_{(i\cap j)^{c}}\bigr{)}.$
We write
$\phi_{i}(\theta)=\phi_{i\cap j}\bigl{(}\theta_{i\cap j}\bigr{)}\,\phi_{i\setminus j}\bigl{(}\theta_{i\setminus j}\bigr{)},\quad\phi_{j}(\theta)=\phi_{i\cap j}\bigl{(}\theta_{i\cap j}\bigr{)}\,\phi_{j\setminus i}\bigl{(}\theta_{j\setminus i}\bigr{)},$
where $\phi_{i\cap j}(\cdot)$ is the maximal monomial factoring both $\phi_{i}(\cdot)$ and $\phi_{j}(\cdot)$, and $\phi_{i\setminus j}$ (resp. $\phi_{j\setminus i}$) is the unique monomial such that
$\phi_{i}(\cdot)=\phi_{i\cap j}(\cdot)\,\phi_{i\setminus j}(\cdot),\qquad\phi_{j}(\cdot)=\phi_{i\cap j}(\cdot)\,\phi_{j\setminus i}(\cdot).$
Then one has:
[…]
and
$\partial^{2}\phi_{i}(\theta)$ $=\begin{pmatrix}\partial^{2}\phi_{i\cap j}\phi_{i\backslash j}&\nabla\phi_{i\cap j}\nabla\phi_{i\backslash j}^{\top}&0&0\\ \nabla\phi_{i\backslash j}\nabla\phi_{i\cap j}^{\top}&\partial^{2}\phi_{i\backslash j}\phi_{i\cap j}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$
Thus
$\partial^{2}\phi_{i}(\theta)\nabla\phi_{j}(\theta)$ $=\begin{pmatrix}\partial^{2}\phi_{i\cap j}\phi_{i\backslash j}&\nabla\phi_{i\cap j}\nabla\phi_{i\backslash j}^{\top}&0&0\\ \nabla\phi_{i\backslash j}\nabla\phi_{i\cap j}^{\top}&\partial^{2}\phi_{i\backslash j}\phi_{i\cap j}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}\begin{pmatrix}\nabla\phi_{i\cap j}\phi_{j\backslash i}\\ 0\\ \phi_{i\cap j}\nabla\phi_{j\backslash i}\\ 0\end{pmatrix}$
$=\begin{pmatrix}\partial^{2}\phi_{i\cap j}\nabla\phi_{i\cap j}\phi_{j\backslash i}\phi_{i\backslash j}\\ \nabla\phi_{i\backslash j}\|\nabla\phi_{i\cap j}\|^{2}\phi_{j\backslash i}\\ 0\\ 0\end{pmatrix}$
$=\phi_{j\backslash i}\begin{pmatrix}\partial^{2}\phi_{i\cap j}\nabla\phi_{i\cap j}\phi_{i\backslash j}\\ \nabla\phi_{i\backslash j}\|\nabla\phi_{i\cap j}\|^{2}\\ 0\\ 0\end{pmatrix}$,
and similarly one has:
$\partial^{2}\phi_{j}(\theta)\nabla\phi_{i}(\theta)$ $=\phi_{i\backslash j}\begin{pmatrix}\partial^{2}\phi_{i\cap j}\nabla\phi_{i\cap j}\phi_{j\backslash i}\\ 0\\ \nabla\phi_{j\backslash i}\|\nabla\phi_{i\cap j}\|^{2}\\ 0\end{pmatrix}$
Finally one has:
[…]
$=\|\nabla\phi_{i\cap j}\|^{2}\begin{pmatrix}0\\ -\nabla\phi_{i\backslash j}\phi_{j\backslash i}\\ \nabla\phi_{j\backslash i}\phi_{i\backslash j}\\ 0\end{pmatrix}$
But as:
$\phi_{j\backslash i}\nabla\phi_{i}-\phi_{i\backslash j}\nabla\phi_{j}=\phi_{i\cap j}\begin{pmatrix}0\\ \phi_{j\backslash i}\nabla\phi_{i\cap j}\\ -\phi_{i\backslash j}\nabla\phi_{j\cap i}\\ 0\end{pmatrix}$
As $\phi_{i\cap j}\neq 0$ (indeed $\theta\in(\mathbb{R}\setminus\{0\})^{D}$, one then has $\begin{pmatrix}0\\ \phi_{j\backslash i}\nabla\phi_{i\cap j}\\ -\phi_{i\backslash j}\nabla\phi_{j\cap i}\end{pmatrix}\in\mathbb{W}_{\phi}(\theta)$ and thus
$\nabla\phi_{i},\nabla\phi_{j}](\theta)\in\mathbb{W}_{\phi}(\theta).$
Appendix F Proof of Corollary˜3.9
We first prove the following proposition (recall that $\mathbb{W}_{\phi}$ is defined in (9) and below).
Proposition F.1.
If $\phi:\mathbb{R}^{D}\to\mathbb{R}^{d}$ is polynomial, then there exists a dense open set $\Theta$ of $\mathbb{R}^{D}$ such that for any $\theta\in\Theta$, $\mathrm{dim}\mathbb{W}_{\phi}(\theta)=\max_{\theta^{\prime}\in\mathbb{R}^{D}}\mathrm{dim}\mathbb{W}_{\phi}(\theta^{\prime})$.
Proof.
Denote $M\coloneqq\max_{\theta^{\prime}\in\mathbb{R}^{D}}\mathrm{dim}\mathbb{W}_{\phi}(\theta^{\prime})\in\mathbb{N}$. Considering $\theta_{0}\in\mathbb{R}^{D}$ such that $\mathrm{dim}\mathbb{W}_{\phi}(\theta_{0})=M$, there exists distinct indices $i_{1},\cdots,i_{M}$ such that the vectors $\nabla\phi_{i_{j}}(\theta_{0})\in\mathbb{R}^{D},1\leq j\leq M$ are linearly independent. There also exists a set $I$ of $M$ coordinates such that the restriction of these vectors to $I$ remains linearly independent. The function $\theta\mapsto\eta(\theta)\coloneqq\text{det}[(\nabla\phi_{i_{1}}(\theta))_{I},\cdots,(\nabla\phi_{i_{M}}(\theta)_{I}]$ is a polynomial on $\mathbb{R}^{D}$ with $\eta(\theta_{0})\neq 0$, hence the set $\mathcal{Z}$ of its zeros is a closed negligeable set of $\mathbb{R}^{D}$. Thus the open dense subset of $\mathbb{R}^{D}$ defined by $\Theta\coloneqq\mathbb{R}^{D}\setminus\mathcal{Z}$ satisfies: for all $\theta\in\Theta$: $\mathrm{dim}\mathbb{W}_{\phi}(\theta)=M$. ∎
See 3.9
Proof.
Since $\phi_{\mathtt{ReLU}}$ is polynomial, we can apply Proposition˜F.1 to obtain an open dense set $\Theta$ on which the dimension of the trace of $\mathbb{W}_{\phi}$ remains constant. By Theorem˜3.8, $\phi_{\mathtt{ReLU}}$ satisfies the Frobenius property. By Proposition˜3.7 every $\theta_{0}\in\Theta$ admits a neighborhood $U$ on which it satisfies the intrinsic recoverability property with respect ot $\phi_{\mathtt{ReLU}}$. By Lemma˜3.2 such a parameter $\theta_{0}$ also satisfies the intrinsic metric property on $U$ with respect to $\phi_{\mathtt{ReLU}}$. ∎
Appendix G Proof of Proposition˜3.10
Lemma G.1.
Let $Y\in\mathbb{R}_{>0}^{n\times m}$. Then there exists a unique pair $(\boldsymbol{\alpha},\boldsymbol{\beta})=:\Gamma(Y)$ of vectors $\boldsymbol{\alpha}\in\mathbb{R}_{>0}^{n}$, $\boldsymbol{\beta}\in\mathbb{R}_{>0}^{m}$ such that
$\boldsymbol{\alpha}^{2}=Y\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m},\quad\text{and}\quad{\boldsymbol{\beta}}^{2}=Y^{\top}\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}.$
Proof.
Define the mappings
$S(\boldsymbol{\beta}):=\sqrt{Y\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m}},\quad T(\boldsymbol{\alpha}):=\sqrt{Y^{\top}\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}}.$
Let $D(a,a^{\prime}):=\|\log(a/a^{\prime})\|_{\infty}$ denote the Thompson metric on $(\mathbb{R}_{+}^{*})^{d}$, where $\mathbb{R}_{+}^{*}$ is the set of positive real numbers. It is known that $\left((\mathbb{R}_{+}^{*})^{d},D\right)$ is a complete metric space. The linear operator $Y$ is 1-Lipschitz with respect to $D$, according to the Birkhoff contraction theorem. Moreover, the square root function is $\frac{1}{2}$-Lipschitz in this metric. Hence, the composition $S\circ T$ is $\frac{1}{4}$-contracting. By the Banach fixed-point theorem, there exists a unique fixed point of $S\circ T$, which implies the existence and uniqueness of the pair $(\boldsymbol{\alpha},\boldsymbol{\beta})$ solving the original equations. ∎
See 3.10
Proof.
Given the general definition of $\phi_{\mathtt{ReLU}}$ (see e.g. Neyshabur et al. (2015); Stock & Gribonval (2022); Gonon et al. (2023)), we study the factorization map
$\phi(u,V,w)\;:=\mathrm{diag}(u)\;V\;\mathrm{diag}(w),$
where $u\in\mathbb{R}^{n},\;V\in\mathbb{R}^{n\times m},\;w\in\mathbb{R}^{m}$ with $u_{i},w_{j}\neq 0$.
Step 1: Gradient flow in parameters.
Let $f:\mathbb{R}^{n\times m}\to\mathbb{R}$ and define the loss $\ell(u,V,w)=f\bigl{(}\phi(u,V,w)\bigr{)}$. Writing $Z=\phi(u,V,w)$ and its gradient $G=\nabla f(Z)$, the gradient-flow ODE (1) $\dot{u}=-\partial_{u}\ell,\;\dot{V}=-\partial_{V}\ell,\;\dot{w}=-\partial_{w}\ell$ is:
$\dot{u}$ $=-\mathrm{Diag}\bigl{(}G\,\mathrm{diag}(w)\,V^{\top}\bigr{)},$
$\dot{V}$ $=-\mathrm{diag}(u)\;G\;\mathrm{diag}(w),$
$\dot{w}$ $=-\mathrm{Diag}\bigl{(}V^{\top}\,\mathrm{diag}(u)\,G\bigr{)},$
Step 2: Induced flow on $z$.
Since $Z=\mathrm{diag}(u)V\mathrm{diag}(w)$, we have
$\dot{Z}=\mathrm{diag}(\dot{u})V\mathrm{diag}(w)+\mathrm{diag}(u)\dot{V}\mathrm{diag}(w)+\mathrm{diag}(u)V\mathrm{diag}(\dot{w}).$
Substituting the above yields
$\dot{Z}=-\mathrm{ddiag}\bigl{(}G\,\mathrm{diag}(w)V^{\top}\bigr{)}V\mathrm{diag}(w)-\mathrm{diag}(u^{2})\,G\,\mathrm{diag}(w^{2})-\mathrm{diag}(u)V\mathrm{ddiag}\bigl{(}V^{\top}\mathrm{diag}(u)G\bigr{)},$
where we set $\mathrm{ddiag}(M)=\mathrm{diag}\bigl{(}\mathrm{Diag}(M)\bigr{)}$.
Eliminating $V$ via $V=\mathrm{diag}(u)^{-1}Z\mathrm{diag}(w)^{-1}$ (possible as $u_{i},w_{j}\neq 0$ on $\Theta$) and using $\mathrm{ddiag}(M\,\mathrm{diag}(a))=\mathrm{ddiag}(M)\,\mathrm{diag}(a)$ one obtains
$\dot{Z}=-\mathrm{ddiag}(Gz^{\top})\,\mathrm{diag}(u^{-2})\,Z-\mathrm{diag}(u^{2})\,G\,\mathrm{diag}(w^{2})-Z\,\mathrm{diag}(w^{-2})\,\mathrm{ddiag}(Z^{\top}G).$
Moreover by Corollary˜3.9 there exists conservation laws $\mathbf{h}$ and a function $\Gamma$ such that $\theta=(u,V,w)=\Gamma(\phi(\theta),\mathbf{h}(\theta))=\Gamma(Z,\mathbf{h}(\theta))$ so that $\boldsymbol{\alpha}:=u^{2}$ and $\boldsymbol{\beta}:=w^{2}$ (entrywise multiplication) can both be expressed as functions $\boldsymbol{\alpha}(Z,\mathbf{h}(\theta))$ and $\boldsymbol{\beta}(Z,\mathbf{h}(\theta))$. Below we explicit such conservation laws and characterize properties of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$.
Step 3: Conserved quantities and elimination of $\boldsymbol{\alpha},\boldsymbol{\beta}$.
The flow (1) preserves the following $n+m$ conservation laws:
$\forall i=1,\dots,n:\quad u_{i}^{2}-\sum_{j=1}^{m}V_{ij}^{2}=\lambda_{i},$
$\forall j=1,\dots,m:\quad w_{j}^{2}-\sum_{i=1}^{n}V_{ij}^{2}=\mu_{j},$
for given constants $\boldsymbol{\lambda}\in\mathbb{R}^{n}$ and $\boldsymbol{\mu}\in\mathbb{R}^{m}$ determined by $\theta_{0}$. Since $V_{ij}=Z_{ij}/(u_{i}w_{j})$, then $(u^{2},w^{2})>0$ is a solution of the coupled system
$u^{2}:~~u_{i}^{4}-\sum_{j=1}^{m}\frac{Z_{ij}^{2}}{w_{j}^{2}}-\lambda_{i}\,u_{i}^{2}$ $=0,$
$w^{2}:~~w_{j}^{4}-\sum_{i=1}^{n}\frac{Z_{ij}^{2}}{u_{i}^{2}}-\mu_{j}\,w_{j}^{2}$ $=0.$
In vector-matrix form (with entrywise squaring):
$\boldsymbol{\alpha}=u^{2},\quad\boldsymbol{\beta}=w^{2},$
$\boldsymbol{\alpha}^{2}-|Z|^{2}\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m}-\boldsymbol{\lambda}\odot\boldsymbol{\alpha}=0,\quad\boldsymbol{\beta}^{2}-(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}-\boldsymbol{\mu}\odot\boldsymbol{\beta}=0,$
where $|Z|^{2}$ is the elementwise square of $Z$ and $\odot$ is the element-wise product.
Special case $\boldsymbol{\lambda}=0,\boldsymbol{\mu}=0$.
Then the system reduces to
$\boldsymbol{\alpha}^{2}=(|Z|^{2})\,\mathrm{diag}(\boldsymbol{\beta})^{-1}\mathbf{1}_{m},\quad\boldsymbol{\beta}^{2}=(|Z|^{2})^{\top}\,\mathrm{diag}(\boldsymbol{\alpha})^{-1}\mathbf{1}_{n}.$ (19)
By Lemma˜G.1 with $Y=|Z|^{2}$ (possible as $Z_{ij}=u_{i}V_{ij}w_{j}\neq 0$ since $\theta\in\Theta$), the exists a unique solution $(\boldsymbol{\alpha},\boldsymbol{\beta})>0$ of the system (19).
In the scalar case $(n=m=1)$ with $|Z|^{2}=z^{2}$ a scalar, the solution is $\boldsymbol{\alpha}=\boldsymbol{\beta}=(|Z|^{2})^{1/3}=|z|^{2/3}.$
Appendix H Proof of Theorem˜4.3.
See 4.3
Proof.
Step 1: rank of $Z$.
As $r\leq\text{min}(n,m)$ and as both $U_{t=0}\in\mathbb{R}^{n\times r}\quad$ and $V_{t=0}\in\mathbb{R}^{m\times r}$ have full rank equal to $r$, it remains the case in a neighborhood $\Omega$ of $\theta_{0}\coloneqq(U_{t=0},V_{t=0})$, and it is also the case for $Z=UV^{\top}$.
Step 2: A quadratic equation for $P:=UU^{\top}$.
Compute
$ZZ^{\top}=UV^{\top}VU^{\top}=U\bigl{(}V^{\top}V\bigr{)}U^{\top}.$
With the hypothesis $U^{\top}U-V^{\top}V=\lambda\mathrm{Id}_{r}$ we get $V^{\top}V=U^{\top}U-\lambda\mathrm{Id}_{r}$, hence
$ZZ^{\top}=U\bigl{(}U^{\top}U-\lambda\mathrm{Id}_{r}\bigr{)}U^{\top}=U\!U^{\top}U\!U^{\top}\;-\;\lambda UU^{\top}=P^{2}-\lambda P.$
Thus $P$ satisfies the quadratic matrix equation
$P^{2}-\lambda P-ZZ^{\top}\;=\;0.$ (20)
Step 3: Simultaneous diagonalisation and scalar reduction.
Write $Z^{\prime}:=ZZ^{\top}$. Because
$P=UU^{\top},\qquad Z^{\prime}=U(V^{\top}V)U^{\top},$
and $U^{\top}U$ differs from $V^{\top}V$ only by a scalar multiple of the identity, we have $(U^{\top}U)(V^{\top}V)=(V^{\top}V)(U^{\top}U)$. Encapsulating by $U$ and $U^{\top}$ yields $PZ^{\prime}=Z^{\prime}P$. Hence $P$ and $Z^{\prime}$ are simultaneously diagonalisable: there exists an orthogonal matrix $W\in\mathbb{R}^{n\times n}$ such that
$P=W\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})W^{\top},\qquad Z^{\prime}=W\operatorname{diag}(\mu_{1},\dots,\mu_{n})W^{\top},$
with $\sigma_{i},\mu_{i}\geq 0$ and where we assume $\sigma_{1}\geq\cdots\geq\sigma_{n}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$.
In the common eigenbasis, (20) becomes for every $i$
$\sigma_{i}^{2}-\lambda\sigma_{i}-\mu_{i}=0.$
Its two roots are
$\sigma_{i}^{\pm}=\frac{\lambda\pm\sqrt{\lambda^{2}+4\mu_{i}}}{2}.$
By the first step, one already has that on $\Omega$, for any $i>r$: $\sigma_{i}=\mu_{i}=0$ so that $\sigma_{i}=\sigma_{i}^{-}$, and that for any $i\leq r$, $\sigma_{i}>0$ and $\mu_{i}>0$. Thus $\sqrt{\lambda^{2}+4\mu_{i}}>|\lambda|$, the “-” root is negative, while $P=UU^{\top}$ is positive-semidefinite. Therefore $\sigma_{i}=\sigma_{i}^{+}$ for $i\leq r$. Let us define $\Pi_{ZZ^{\top}}\coloneqq W\operatorname{diag}(\underbrace{1,\cdots,1}_{\times r},0,\cdots,0)W^{\top}$ the orthogonal projector on $\mathrm{range}(ZZ^{\top})$. It follows that:
$P=\Pi_{ZZ^{\top}}\times\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right].$ (21)
Step 4: The expression for $Q:=VV^{\top}$. A fully analogous computation gives
$Z^{\top}Z=VU^{\top}UV^{\top}=V\bigl{(}V^{\top}V+\lambda\mathrm{Id}_{r}\bigr{)}V^{\top}=Q^{2}+\lambda Q,$
so that $Q$ satisfies
$Q^{2}+\lambda Q-Z^{\top}Z\;=\;0.$ (22)
Because $Q$ and $T:=Z^{\top}Z$ commute, they share an orthonormal eigenbasis in which (22) reduces to
$\tau_{i}^{2}+\lambda\tau_{i}-\mu_{i}=0\quad(\tau_{i}\!\geq\!0,\;\mu_{i}\!\geq\!0).$
By the first step, one already has that on $\Omega$, for any $i>r$: $\tau_{i}=\mu_{i}=0$ and that for any $i\leq r$, $\tau_{i}\neq 0$ and $\mu_{i}\neq 0$. For $i\leq r$ the positive root (as $\sqrt{\lambda^{2}+4\mu_{i}}>|\lambda|$) is
$\tau_{i}=\frac{-\lambda+\sqrt{\lambda^{2}+4\mu_{i}}}{2},$
so that
$Q=\Pi_{Z^{\top}Z}\times\left[-\frac{\lambda}{2}I_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,T}\right],$ (23)
with $T=Z^{\top}Z$ and where $\Pi_{Z^{\top}Z}$ is the orthogonal projector on $\mathrm{range}(Z^{\top}Z)$.
Step 5: Uniqueness and conclusion In both cases (21)–(23) are the only solutions consistent with $UU^{\top}\succeq 0$ and $VV^{\top}\succeq 0$ and with $\mathrm{rank}(Z)=r$ on $\Omega$. Finally one has on $\Omega$:
$\overset{.}{Z}$ $=-UU^{\top}\nabla f(Z)-\nabla f(Z)VV^{\top}$
$=-\Pi_{ZZ^{\top}}\times\left[\frac{\lambda}{2}\mathrm{Id}_{n}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{n}+4\,ZZ^{\top}}\right]\nabla f(X)-\nabla f(X)\Pi_{Z^{\top}Z}\times\left[-\frac{\lambda}{2}I_{m}+\frac{1}{2}\sqrt{\lambda^{2}\mathrm{Id}_{m}+4\,Z^{\top}Z}\right],$
which concludes the proof. ∎
Appendix I Proof of Theorem˜4.4
We first show the following lemma:
Lemma I.1.
If $S\neq\lambda\mathrm{Id}_{r}$ with $S$ a real symmetric matrix, then there exists a skew-symmetric matrix $A$ such that $A,S]\neq 0$.
Proof.
Let us assume $S\neq\lambda\mathrm{Id}_{r}$ (in particular $r>1$ necessarily). Thus there are at least two distinct eigenvalues of $S$ $\delta$ and $\mu$ associated to the eigenvectors $x$ and $y$. Then $A\coloneqq xy^{\top}-yx^{\top}\neq 0$ is a skew-symmetric matrix that satisfies:
$[A,S]$ $=(xy^{\top}-yx^{\top})S-S(xy^{\top}-yx^{\top})$
$=x(Sy)^{\top}-y(Sx)^{\top}-(Sx)y^{\top}+(Sy)x^{\top}\text{ as $S$ is symmetric}$
$=\mu xy^{\top}-\delta yx^{\top}-\delta xy^{\top}+\mu yx^{\top}$
$=(\underbrace{\mu-\delta}_{\neq 0})(xy^{\top}+yx^{\top})\neq 0,$
as $\mu\neq\delta$, and which concludes the proof. ∎
See 4.4
Proof.
In light of the necessary condition of Theorem˜2.14 we will first characterize $\mathrm{ker}\partial M(\theta)$ for any $\theta=(U,V)$. Then, with $\mathbf{h}(\theta)=U^{\top}U-V^{\top}V$ and $\phi(\theta)=\phi_{\mathtt{Lin}}(\theta)=UV^{\top}$, we will exhibit a subspace $\mathcal{V}$ of $\mathrm{ker}\partial\mathbf{h}(\theta)\cap\mathrm{ker}\partial\phi(\theta)$ such that $\mathcal{V}\subsetneq\mathrm{ker}\partial M(\theta)$. We will then conclude using the needed calculus and  Theorem˜2.14.
Step 1: Characterization of $\mathrm{ker}\partial M(\theta)$ for any $\theta=(U,V)$.
By (11) (with $L=2$, $U_{2}=U,U_{1}=V^{\top}$), one can write $M(\theta)\mathrm{vec}(X)=\mathrm{vec}(UU^{T}X+XVV^{\top})$ for any matrix $X\in\mathbb{R}^{n\times m}$. Using the Kronecker product and the fact that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$, this expression can be rewritten as:
$M(\theta)=\mathrm{Id}_{m}\otimes(UU^{T})+(VV^{\top})\otimes\mathrm{Id}_{n}.$ (24)
Thus differentiating (24) yields that for any $(H,K)$ of the same dimensions as $(U,V)$ we have $\partial M(\theta).(H,K)=\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})+{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$, and thus: $(H,K)\in\mathrm{ker}\partial M(\theta)$ if and only if $\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}$. We now show that
$\mathrm{ker}\partial M(\theta)=\left\{(H,K):\exists\mu\in\mathbb{R},UH^{\top}+HU^{\top}=\mu\mathrm{Id}_{n}\,\text{and}\,VK^{\top}+KV^{\top}=-\mu\mathrm{Id}_{m}\right\}.$ (25)
The converse inclusion is clear. We now prove the direct inclusion. Let us consider $(H,K)\in\mathrm{ker}\partial{M}(\theta)$, then one has $\mathrm{Id}_{m}\otimes(UH^{T}+HU^{T})=-{(VK^{\top}+KV^{\top})}\otimes\mathrm{Id}_{n}.$ Still using that $(A\otimes B)\mathrm{vec}(X)=\mathrm{vec}(BXA^{\top})$ and denoting $U^{\prime}\coloneqq UH^{T}+HU^{T}$ and $V^{\prime}\coloneqq{(VK^{\top}+KV^{\top})}$, this exactly means that for any matrix $X\in\mathbb{R}^{n\times m}$ one has $U^{\prime}X=-XV^{\prime\top}$. To conclude, we only need to show that this implies the existence of $\mu,{\mu^{\prime}}\in\mathbb{R}$ such that $U^{\prime}=\mu\mathrm{Id}_{n}\,\text{and}\,V^{\prime}=-{\mu^{\prime}}\mathrm{Id}_{m}$, since the equality $U^{\prime}X=-XV^{\prime\top}$ then also implies $\mu=\mu^{\prime}$. This is immediate if $V^{\prime}=0$ since in this case $U^{\prime}$ must also be equal to zero as $U^{\prime}X=0$ for every $X$. Assume now that $V^{\prime}$ is non-zero so there exists a vector $v$ such that $V^{\prime\top}v\neq 0$. Considering any such $v$ and any vector $u$, and setting $X=uv^{\top}$, we have
$(U^{\prime}u)v^{\top}=U^{\prime}X=-XV^{\prime}=-u(V^{\prime\top}v)^{\top}$
hence $U^{\prime}u$ is colinear with $u$. Since this holds for any choice of $u$, we deduce indeed that $U^{\prime}$ is proportional to $\mathrm{Id}_{n}$. A similar reasoning yields that $V^{\prime}\propto\mathrm{Id}_{m}$. This concludes the proof of (25).
Step 2: Characterization of a subspace $\mathcal{V}\subseteq\mathrm{ker}\partial\mathbf{h}(\theta)\cap\mathrm{ker}\partial\phi(\theta)$. Since $\mathbf{h}(\theta)=U^{\top}U-V^{\top}V$ and $\phi(\theta)=UV^{\top}$ we have
$\partial h(\theta).(H,K)$ $=U^{\top}H+H^{\top}U-V^{\top}K-K^{\top}V$
$\partial\phi(\theta).(H,K)$ $=UK^{\top}+HV^{\top}$
and one can easily check that for any $\theta$ such that $\mathbf{h}(\theta)=S$ we have
$\mathcal{V}$ $:=\left\{\begin{pmatrix}U\Delta\\ -V\Delta^{\top}\end{pmatrix}:\Delta\in\mathbb{R}^{r\times r},(\Delta^{\top}+\Delta)\,U^{\top}U+U^{\top}U\,(\Delta+\Delta^{\top})=\Delta\,S+S\,\Delta^{\top}\right\}$
$\subseteq\mathrm{ker}\partial\mathbf{h}(\theta)\cap\mathrm{ker}\partial\phi(\theta).$
Step 3: Proof that $\mathcal{V}\not\subset\mathrm{ker}\partial M(\theta)$. The fact that a matrix $\Delta\in\mathbb{R}^{r\times r}$ satisfies
$(\Delta^{\top}+\Delta)\,U^{\top}U+U^{\top}U\,(\Delta+\Delta^{\top})=\Delta\,S+S\,\Delta^{\top},$
is equivalent to
$\Delta_{S}(2U^{\top}U-S)+(2U^{\top}U-S)\Delta_{S}=[\Delta_{A},S].$
with $\Delta_{S}$ (resp. $\Delta_{A}$) the symmetric (resp. skew symmetric) part of $\Delta$ (so that $\Delta=\Delta_{S}+\Delta_{A}$). Denote $\mathcal{S}_{r}$ (resp. $\mathcal{A}_{r}$) the set of $r\times r$ symmetric (resp. skew symmetric) matrices and $L$ the Lyapunov operator defined by:
$L:\Delta_{S}\in\mathcal{S}_{r}\mapsto L(\Delta_{S})$ $\coloneqq\Delta_{S}(2U^{\top}U-S)+(2U^{\top}U-S)\Delta_{S}$
$=\Delta_{S}(U^{\top}U+V^{\top}V)+(U^{\top}U+V^{\top}V)\Delta_{S}\in\mathcal{S}_{r}$
We obtain
$\mathcal{V}$ $=\left\{\begin{pmatrix}U(\Delta_{S}+\Delta_{A})\\ -V(\Delta_{S}-\Delta_{A})\end{pmatrix}:(\Delta_{S},\Delta_{A})\in\mathcal{S}_{r}\times\mathcal{A}_{r},L(\Delta_{S})=[\Delta_{A},S]\right\}.$
As $S\neq\lambda\mathrm{Id}_{r}$, by Lemma˜I.1 there exists a skew-symmetric matrix $\Delta_{A}\in\mathcal{A}_{r}$ such that $\Delta_{A},S]\neq 0$. As $U^{\top}U+V^{\top}V$ is positive definite (as either $U$ or $V$ has full column-rank) its eigenvalues $\lambda_{i}>0$ satisfy $\lambda_{i}+\lambda_{j}\neq 0$, so (see e.g. Bartels & Stewart (1972)) in particular the Lyapunov operator: $L:\mathcal{S}_{r}\to\mathcal{S}_{r}$ is invertible. Since $\Delta_{A},S]=\Delta_{A}S-S\Delta_{A}\in\mathcal{S}_{r}$, we obtain that there exists $\Delta_{S}\neq 0$ such that $L(\Delta_{S})=[\Delta_{A},S]$. This particular choice of $\Delta_{S}$ and $\Delta_{A}$ exhibits a parameter $\theta^{\prime}=(U\Delta,-V\Delta^{\top})$ that satifies $\theta^{\prime}\in\mathcal{V}\subseteq\mathrm{ker}\partial\phi(\theta)\cap\mathrm{ker}\partial\mathbf{h}(\theta)$. We now show that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$. We proceed by contradiction: if $\theta^{\prime}\in\mathrm{ker}\partial{M}(\theta)$ then, by (25), there exists $\mu\in\mathbb{R}$ such that $U(\Delta^{\top}+\Delta)U^{\top}=\mu\mathrm{Id}_{n}\quad\text{and}\quad V(\Delta^{\top}+\Delta)V^{\top}=-\mu\mathrm{Id}_{m}$ that is to say
$2U\Delta_{S}U^{\top}=\mu\mathrm{Id}_{n}\quad\text{and}\quad 2V\Delta_{S}V^{\top}=-\mu\mathrm{Id}_{m}.$ (26)
When $r\leq\max(m,n)$ and since $U$, $V$ are full rank, at least one of the two matrices $U$ or $V$ is full column rank $r$. Without loss of generality, let us assume that $U$ is full column rank. Then $U^{\top}U$ is invertible and we deduce that,
$2\Delta_{S}=\mu(U^{\top}U)^{-1}.$ (27)
Moreover if (as we indeed show below) $\mathrm{range}U^{\top}\cap\mathrm{range}V^{\top}\neq\{0\}$, then by considering $z=U^{\top}x=V^{\top}y\neq 0$ for some $x,y\in\mathbb{R}^{r}$, one deduces from (26) that $\mu\|x\|_{2}^{2}=2x^{\top}U\Delta_{S}U^{\top}x=z^{\top}\Delta_{S}z=2y^{\top}V\Delta_{S}V^{\top}y=-\mu\|y\|_{2}^{2}$ and thus $\mu=0$. Hence $\Delta_{S}=0$ by (27), contradicting $L(\Delta_{S})=[\Delta_{A},S]\neq 0$, which shows that $\theta^{\prime}\notin\mathrm{ker}\partial{M}(\theta)$.
Thus we only need to prove that one has $\mathrm{range}U^{\top}\cap\mathrm{range}V^{\top}\neq\{0\}$, and indeed:
$\mathrm{dim}(\mathrm{range}(U^{\top})\cap\mathrm{dim}(\mathrm{range}(V^{\top}))$ $=\underbrace{\mathrm{rank}(U^{\top})}_{=\mathrm{rank}(U)}+\underbrace{\mathrm{rank}(V^{\top})}_{=\mathrm{rank}(V)}-\mathrm{dim}(\underbrace{\mathrm{range}(U^{\top})+\mathrm{range}(V^{\top})}_{\mathrm{range}((U^{\top}|\,V^{\top}))})$
$=\underbrace{\mathrm{rank}(U)+\mathrm{rank}(V)}_{\geq\min(r,n)+\min(r,m)\geq r+1}-\underbrace{\mathrm{rank}\left(\begin{pmatrix}U\\ V\end{pmatrix}\right)}_{=r}>0,$
where we used in the last line that $r\leq\text{max}(n,m)$.
Step 4: Conclusion.
As both $U_{t=0}$ and $V_{t=0}$ have full rank it remains the case in a neighborhood $\Omega$ of $\theta_{0}$. Moreover as $r\leq\text{max}(n,m)$ then one of the two matrices has a full column rank on $\Omega$. In particular the vertical concatenation $\begin{pmatrix}U\\ V\end{pmatrix}$ has full rank (equal to $r$) on $\Omega$ as $r\leq\text{max}(n,m)\leq n+m$.
Since $(\begin{smallmatrix}U\\ V\end{smallmatrix})$ has full rank on $\Omega$, by (Marcotte et al., 2023, Proposition 4.2 and Corollary 4.4) the vector-valued function $\mathbf{h}$ contains a complete set of conservation laws.
We now show by contradiction that for any ${\Omega^{\prime}}\subseteq{\Omega}$, $\theta_{0}$ does not satisfy the intrinsic metric on $\Omega^{\prime}$. Let us assume there exists a neighborhood $\Omega^{\prime}\subseteq\Omega$ of $\theta_{0}$ and a set of conservation laws $\mathbf{h}_{0}$ for $\phi$ and a function $K_{\theta_{0}}$ such that $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ for each $\theta\in\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\cap\Omega^{\prime}$, where $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}$. As the family of conservation laws $\mathbf{h}$ is complete on $\Omega$ (and in particular on $\Omega^{\prime}$) and as $\mathrm{Lie}(\mathbb{W}_{\phi})(\theta)$ has a constant dimension on $\Omega$ (and thus on $\Omega^{\prime}$) by (Marcotte et al., 2023, Proposition 4.3), using (Marcotte et al., , Proposition 2.12) yields that $\mathcal{M}_{\theta_{0}}^{\mathbf{h_{0}}}\coloneqq\{\theta:\mathbf{h_{0}}(\theta)=\mathbf{h_{0}}(\theta_{0})\}\supset\mathcal{M}_{\theta_{0}}^{\mathbf{h}}\coloneqq\{\theta:\mathbf{h}(\theta)=\mathbf{h}(\theta_{0})\}$. Thus the function $K_{\theta_{0}}$ also satisfies $M(\theta)=K_{\theta_{0}}(\phi(\theta))$ on $\mathcal{M}_{\theta_{0}}^{\mathbf{h}}$, hence $\mathbf{h}$ satisfies assumption $i)$ of Theorem˜2.14. As the rank of $\partial h(\theta)$ is constant on $\Omega^{\prime}$, we deduce by Theorem˜2.14 the inclusion (7), which contradicts the previous step. ∎
Appendix J Proof of Proposition˜4.5.
See 4.5
Proof.
Since $\partial\phi(\theta)=[v^{\top},u^{\top}]$ we have
$\partial\phi(\theta)\partial\phi(\theta)^{\top}=\|u\|^{2}+\|v\|^{2}.$
Since $\mathbf{h}(\theta):=uu^{\top}-vv^{\top}$ is a conservation law of $\phi_{\mathtt{Lin}}$ for every $\theta=(u,v)$ on the trajectory one has: $S=uu^{\top}-vv^{\top}$, and therefore $S^{2}=\|u\|^{2}uu^{\top}-zuv^{\top}-zvu^{\top}+\|v\|^{2}vv^{\top}$. Thus
$\mathrm{tr}(S^{2})=\|u\|^{4}+\|v\|^{4}-2z^{2}.$
As one also has: $(\|u\|^{2}-\|v\|^{2})^{2}=\mathrm{tr}(S)^{2}$, one has:
$(\partial\phi(\theta)\partial\phi(\theta)^{\top})^{2}=(\|u\|^{2}+\|v\|^{2})^{2}$ $=2(\|u\|^{4}+\|v\|^{4})-(\|u\|^{2}-\|v\|^{2})^{2}$
$=2(\mathrm{tr}(S^{2})+2z^{2})-\mathrm{tr}(S)^{2}$
$=2\mathrm{tr}(S^{2})+4z^{2}-\mathrm{tr}(S)^{2},$
which concludes the proof. ∎
Appendix K Proof of Theorem˜4.6.
See 4.6
Proof.
Let us first outline the main steps of the proof. We first show that the equalities $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$ and $Z_{L}^{\top}Z_{L}=R_{L-1}(U_{1}^{\top}U_{1})$ hold on the whole trajectory. Then we prove that this implies the expression of $S_{j}$ (resp. of $T_{j}$) in terms of $U_{L}U_{L}^{\top}$ (resp. of $U_{1}^{\top}U_{1}$) along the whole trajectory too. Finally we show that along the whole trajectory $U_{L}U_{L}^{\top}$ and $U_{1}^{\top}U_{1}$ (and therefore all $S_{j}$’s and $T_{j}$’s) are entirely characterized by $Z_{L}=\phi_{\mathtt{Lin}}(\theta_{L})$ and the initial conditions (captured by $\boldsymbol{\lambda}$). This will thus imply that $\theta_{L}(0)$ satisfies the intrinsic dynamic property on $\mathbb{R}^{D}$ with respect to $\phi_{\mathtt{Lin}}$.
Step 1: Expression of $Z_{L}Z_{L}^{\top}$ as a polynomial in $U_{L}U_{L}^{\top}$
Since $U_{j+1}^{\top}U_{j+1}-U_{j}U_{j}^{\top}$ is a set of conservation laws for $\phi_{\mathtt{Lin}}$, the fact that the relaxed balanced conditions (13) hold at initialization implies that they hold along the whole trajectory.
We prove by induction on $1\leq\ell\leq L$ that $Z_{\ell}:=U_{\ell}\ldots U_{1}$ satisfies $Z_{\ell}Z_{\ell}^{\top}=P_{\ell}(U_{\ell}U_{\ell}^{\top})$ for some polynomial $P_{\ell}$ of degree $\ell$ that satisfy $P_{1}(x)=x$ and $P_{\ell}(x)=xP_{\ell-1}(x-\lambda_{\ell-1})$ for $2\leq\ell\leq L$. For $\ell=1$ we trivially have $Z_{\ell}=U_{\ell}$ hence the result is true. Now consider $2\leq\ell\leq L$ and assume that the result holds true for $\ell-1$. Since $Z_{\ell}=U_{\ell}Z_{\ell-1}$ we have
[…]
where we used (13) for $i=\ell-1$. Denoting $\hat{P}_{\ell-1}(x):=P_{\ell-1}(x-\lambda_{\ell-1})$ we obtain $Z_{\ell}Z_{\ell}^{\top}=U_{\ell}\hat{P}_{\ell-1}(U_{\ell}^{\top}U_{\ell})U_{\ell}^{\top}=U_{\ell}U_{\ell}^{\top}\hat{P}_{\ell-1}(U_{\ell}U_{\ell}^{\top})=P_{\ell}(U_{\ell}U_{\ell}^{\top})$. This concludes the induction.
Given the recursion formula for $P_{\ell}$, another easy induction yields
$P_{\ell}(x)=\Pi_{k=0}^{\ell-1}(x-\sum_{i=1}^{k}\lambda_{\ell-i}),\quad 1\leq\ell\leq L.$ (28)
Specializing to $\ell=L$ we obtain $P_{L}=Q_{0}$ as claimed.
Step 2: Expression of $S_{j}$ (resp. of $Z_{L}^{\top}Z_{L}$ and $T_{j}$) as a polynomial in $U_{L}U_{L}^{\top}$ (resp. in $U_{1}^{\top}U_{1}$).
It is a direct consequence of the first step, as we now explain. To show the result on $S_{j}$, consider the new variable $\theta^{\prime}=(U^{\prime}_{L-j},\ldots,U^{\prime}_{1}):=(U_{L},\ldots,U_{j+1})$ and $Z^{\prime}\coloneqq U^{\prime}_{L-j}\cdots U^{\prime}_{1}=U_{L}\cdots U_{j+1}$. With these notations we have $S_{j}=Z^{\prime}Z^{\prime\top}$, and the relaxed balanced conditions imply that:
$(U^{\prime}_{i+1})^{\top}U^{\prime}_{i+1}-U^{\prime}_{i}(U^{\prime}_{i})^{\top}=\lambda^{\prime}_{i}\mathrm{Id}_{n},\quad 1\leq i\leq L-j-1$
where $\boldsymbol{\lambda}^{\prime}=(\lambda^{\prime}_{L-j-1},\ldots,\lambda^{\prime}_{1}):=(\lambda_{L-1},\ldots,\lambda_{j+1})$. By the first step, we obtain the desired expression.
Similar computations with $\theta^{\prime}=(U_{1}^{\top},\ldots,U_{j-1}^{\top})$, $Z^{\prime}=U_{1}^{\top}\ldots U_{j-1}^{\top}$ and ${\boldsymbol{\lambda}}^{\prime}=(-\lambda_{1},\ldots,-\lambda_{j-2})$ show the desired expression for $T_{j}=Z^{\prime}Z^{\prime\top}$ and $Z_{L}^{\top}Z_{L}$ as well.
Step 3: Characterization of $U_{L}U_{L}^{\top}$ via $Z_{L}$ and the initial conditions. The proof that $U_{1}^{\top}U_{1}$ is characterized by $Z_{L}$ (in fact $Z_{L}^{\top}Z_{L}$) and the initial conditions is similar and therefore omitted.
By the first step we have $Z_{L}Z_{L}^{\top}=Q_{0}(U_{L}U_{L}^{\top})$, hence $U_{L}U_{L}^{\top}$ is indeed a matrix root of this equation. As both matrices $Z_{L}Z_{L}^{\top}$ and $U_{L}U_{L}^{\top}$ are real symmetric, the above expression shows that we can reduce to the scalar study of their eigenvalues.
As we show below, a consequence of the relaxed balancedness conditions (13) is that all eigenvalues of the positive semi-definite matrix $U_{L}U_{L}^{\top}$ belong to the interval $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Thus, considering any eigenvalue $e\geq 0$ of the positive semi-definite matrix $Z_{L}Z_{L}^{\top}$, it is enough to show that the polynomial equation $R(X):=Q_{0}(X)-e=0$ admits a unique root in this interval.
The existence of a root in $I$ is a consequence of the mean value theorem, since $R(\text{max}(0,a_{1},\cdots,a_{L-1}))=-e\leq 0$ and $\lim_{x\to\infty}R(x)=+\infty$. To prove uniqueness, we proceed by contradiction: assume that $R(X)$ admits two distinct roots $x_{1}<x_{2}$ in $I$. By Rolle’s theorem $R^{\prime}(X)=Q_{0}^{\prime}(X)$ has a root in $x_{1},x_{2}[$. This contradicts the fact that, by the construction of $Q_{0}$ and Rolle’s theorem, all roots of $Q_{0}^{\prime}(X)$ are contained in the open interval $(\text{min}(0,a_{1},...,a_{L-1}),\text{max}(0,a_{1},...,a_{L-1}))$.
To conclude the proof, we show that indeed all eigenvalues of $U_{L}U_{L}^{\top}$ belong to $I:=[\max(0,a_{1},\ldots,a_{L-1}),\infty)$. Denote $\sigma_{i}=\inf\text{sp}(U_{i}U_{i}^{\top})$, $1\leq i\leq L$. Since each matrix $U_{\ell}$ is square and positive semi-definite, we have $\text{sp}(U_{i}U_{i}^{\top})=\text{sp}(U_{i}^{\top}U_{i})\subseteq[0,\infty)$ for every $1\leq i\leq L$, and by (13) we also have $\text{sp}(U_{i+1}U_{i+1}^{\top})=\lambda_{i}+\text{sp}(U_{i}U_{i}^{\top})$, hence $\sigma_{i+1}=\sigma_{i}+\lambda_{i}\geq 0$ for $1\leq i\leq L-1$. An easy recursion shows that $\sigma_{i}\geq\max(0,\sum_{j=1}^{i-1}\lambda_{j})$ for $1\leq i\leq L$, hence the result. ∎
We now anticipate a slight generalization part of the results of Theorem˜4.6 that will be used later in the proof of Theorem˜4.8.
Lemma K.1 (Perturbed relaxed balanced condition).
Consider matrices $(U_{k})_{k=0}^{L-1}\subset\mathbb{R}^{n\times n}$ and scalars $(\lambda_{k})_{k=0}^{L-1}$. Denoting $h\coloneqq 1/L$, define
$C_{U}$ $\coloneqq\max(1,\max_{k}\|U_{k}\|),\quad C_{\lambda}\coloneqq\max_{k}|\lambda_{k}|$ (29)
$\eta$ $\coloneqq L^{2}\cdot\max_{0\leq k\leq L-2}\|(U_{k+1}^{\top}U_{k+1}-U_{k}U_{k}^{\top})-h^{2}\lambda_{k}\,\mathrm{Id}_{n}\|$ (30)
Fix $j\in\{0,\dots,L-2\}$ and recall that $S_{j}:=(U_{L-1}\cdots U_{j+1})(U_{L-1}\cdots U_{j+1})^{\top}.$ Define $a_{0}:=0$ and, for $k\geq 1$, $a_{k}:=h^{2}\sum_{i=1}^{k}\lambda_{L-1-i}$, $C_{0}\coloneqq 2C_{\lambda}$, $C_{1}\coloneqq(C_{U}^{2}+\eta h^{2}-1)/h$. Then
$\max_{j}\|S_{j}-\prod_{k=0}^{L-1-(j+1)}\big{(}U_{L-1}U_{L-1}^{\top}-a_{k}\mathrm{Id}_{n}\big{)}\|\leq\left(C_{0}e^{C_{1}}e^{C_{0}(1+C_{1})}+e^{C_{1}}\right)\eta.$ (31)
Before proving this lemma, we state the following lemma, as it will be essential in the proof of Lemma˜K.1: it provides a uniform bound on the Lipschitz constant of a class of polynomials.
Lemma K.2 (Uniform Lipschitz bound).
Consider $C_{0}>0$, $C_{1}>0$. For any $0<h\leq 1$, any integer $1\leq d\leq 1/h$, any degree–$d$ polynomial
$Q_{d}(x)\;=\;\prod_{k=1}^{d}(x-c_{k}),$
with $\max_{k}|c_{k}|\;\leq\;C_{0}h$, and any matrices $A,\;A+\Delta\in B_{R(h)}:=\{X:\|X\|\leq R(h)\}$ where $R(h)\coloneqq 1+C_{1}h$ and where $\|\cdot\|$ denotes the Frobenius norm, one has
$\|Q_{d}(A+\Delta)-Q_{d}(A)\|\;\leq\;\frac{K}{h}\,\|\Delta\|,\ \text{with}\ K=K(C_{0},C_{1})=C_{0}e^{C_{0}(1+C_{1})}+e^{C_{1}}.$ (32)
Proof.
Step 1: Scalar Lipschitz constant on the ball $B_{R}$. For any matrix polynomial $Q(x)=\sum_{m=0}^{d}\alpha_{m}x^{m}$ one has, denoting $DQ(X)[H]=\sum_{m=1}^{d}\alpha_{m}\sum_{j=0}^{m-1}X^{j}HX^{m-1-j}$:
$Q(A+\Delta)-Q(A)$ $=\int_{0}^{1}\!DQ\!\bigl{(}A+t\Delta\bigr{)}[\Delta]\,dt,$
$\|DQ(X)[H]\|$ $\leq L_{Q}({\|X\|_{2\to 2}})\,\|H\|{\leq L_{Q}(\|X\|)\,\|H\|},\quad\forall X,\ \forall H$
where $L_{Q}(R):=\sum_{m=1}^{d}|\alpha_{m}|\,m\,R^{m-1}$ (we used here that the spectral norm is bounded by the Frobenius norm).
Step 2: Bounding $L_{Q_{d}}(R(h))$. Exploiting the coefficient–root relation on $Q_{d}$ that is unitary yields $|\alpha_{m}|\leq\binom{d}{m}\beta^{m}$ where $\beta\coloneqq C_{0}h$ for any $0\leq m\leq d-1$. Since $\alpha_{d}=1$, for any $R>0$ we obtain
$L_{Q_{d}}(R)\;\leq\;\beta\sum_{m=1}^{d}\binom{d}{m}m(\beta R)^{m-1}+dR^{d-1}=d\beta(1+\beta R)^{d-1}+dR^{d-1}.$
Insert $d-1\leq d\leq 1/h$, $\beta=C_{0}h$. Since $R(h)=1+C_{1}h\leq R(1)=1+C_{1}$ (as $h\leq 1$) we get:
$L_{Q_{d}}(R(h))\;\leq\;\frac{1}{h}\,C_{0}h\,(1+C_{0}hR(h))^{1/h}+d(1+C_{1}h)^{1/h}\leq C_{0}\,e^{C_{0}R(h)}+\frac{e^{C_{1}}}{h}\leq\frac{C_{0}\,e^{C_{0}(1+C_{1})}+e^{C_{1}}}{h}.$
where the exponential bound uses $(1+t)^{1/t}\leq e$ for $t>0$. We define $K=K(C_{0},C_{1})\coloneqq C_{0}e^{C_{0}(1+C_{1})}+e^{C_{1}}$.
Step 3: Conclusion. Applying the integral formula of Step 1with the bound from Step 2 gives
$\|Q_{d}(A+\Delta)-Q_{d}(A)\|\;\leq\;(K/h)\,\|\Delta\|,$
for every $A,\Delta$ with $A,A+\Delta\in B_{R(h)}$, which is (32). ∎
We now prove Lemma˜K.1.
Proof.
Step 0: Reindexing. Work with the truncated sequence $(U^{\prime}_{1},\dots,U^{\prime}_{N}):=(U_{j+1},\dots,U_{L-2},U_{L-1})$, where N:=L1jN:=L-1-j. Define $Z_{\ell}:=U^{\prime}_{\ell}\cdots U^{\prime}_{1}$ for $1\leq\ell\leq N$ and $M_{\ell}:=U^{\prime}_{\ell}U^{\prime}_{\ell}{}^{\top}$. Then $S_{j}=Z_{N}Z_{N}^{\top}$.
We also observe that by the definition of $\eta$ in (30), since $h=1/L$, we have for each $1\leq\ell\leq N$
${U^{\prime}_{\ell}}^{\top}U^{\prime}_{\ell}-M_{\ell-1}={U^{\prime}_{\ell}}^{\top}U^{\prime}_{\ell}-U^{\prime}_{\ell-1}{U^{\prime}_{\ell-1}}^{\top}=h^{2}{\lambda_{\ell+j-1}}\mathrm{Id}_{n}+r^{\prime}_{\ell-1},\qquad\|r^{\prime}_{\ell-1}\|\leq h^{2}\eta.$ (33)
Step 1: Polynomial representation with a perturbation. We prove by induction on $\ell$ that
$E_{\ell}\coloneqq Z_{\ell}Z_{\ell}^{\top}-P_{\ell}(M_{\ell})\quad\text{satisfies}\ \|E_{\ell}\|\leq\ell KC_{U}^{{2\ell}}\cdot h\eta$ (34)
where the polynomials $P_{\ell}$ are defined by
$P_{1}(x)\coloneqq x,\qquad P_{\ell}(x)\coloneqq x\,P_{\ell-1}\bigl{(}x-b_{\ell-1}\bigr{)}\quad(2\leq\ell\leq N),$
with $b_{\ell-1}\coloneqq h^{2}{\lambda_{\ell+j-1}}$ (matching the re-indexed sequence), and the constant $K$ is obtained by Lemma˜K.2 applied to the constants $C_{0}\coloneqq 2C_{\lambda}$ and $C_{1}\coloneqq(C_{U}^{2}+\eta h^{2}-1)/h$.
Base case $\ell=1$. Trivial: $Z_{1}=U^{\prime}_{1}$, so $Z_{1}Z_{1}^{\top}=M_{1}=P_{1}(M_{1})$ and $E_{1}=0$.
Induction step. Assume (34) holds at rank $\ell-1$. Since $Z_{\ell}=U^{\prime}_{\ell}Z_{\ell-1}$,
$Z_{\ell}Z_{\ell}^{\top}=U^{\prime}_{\ell}\bigl{(}Z_{\ell-1}Z_{\ell-1}^{\top}\bigr{)}U^{\prime}_{\ell}{}^{\top}=U^{\prime}_{\ell}P_{\ell-1}(M_{\ell-1})U^{\prime}_{\ell}{}^{\top}\;+\;U^{\prime}_{\ell}E_{\ell-1}U^{\prime}_{\ell}{}^{\top}.$
By induction hypothesis and the fact that the spectral norm is bounded by the Frobenius norm, the second term of the right hand side is bounded as
$\|U_{\ell}^{\prime}E_{\ell-1}U^{\prime}_{\ell}{}^{\top}\|\leq C_{U}^{2}\|E_{\ell-1}\|\leq C_{U}^{2}(\ell-1)KC_{U}^{{2(\ell-1)}}h\eta\leq(\ell-1)C_{U}^{{2\ell}}Kh\eta,$
hence we only need to show that
$\|U^{\prime}_{\ell}P_{\ell-1}(M_{\ell-1})U^{\prime}_{\ell}{}^{\top}-P_{\ell}(M_{\ell})\|\leq C_{U}^{{2\ell}}K\cdot h\eta.$
Write $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$. From (33) and the definition of $M_{\ell-1}=U^{\prime}_{\ell-1}[U^{\prime}_{\ell-1}]^{\top}$ one gets
$M_{\ell-1}=U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-b_{\ell-1}\mathrm{Id}_{n}-r_{\ell-1}^{\prime},\qquad\|r_{\ell-1}^{\prime}\|\leq h^{2}\eta.$
Hence
$U^{\prime}_{\ell}P_{\ell-1}(M_{\ell-1})U^{\prime}_{\ell}{}^{\top}$ $=U^{\prime}_{\ell}Q_{\ell-1}\big{(}U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-r_{\ell-1}^{\prime}\big{)}U^{\prime}_{\ell}{}^{\top}$
$=\underbrace{U^{\prime}_{\ell}Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell})U^{\prime}_{\ell}{}^{\top}}_{=M_{\ell}\,Q_{\ell-1}(M_{\ell})=P_{\ell}(M_{\ell})}+U^{\prime}_{\ell}\!\Big{(}Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-r_{\ell-1}^{\prime})-Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell})\Big{)}\!U^{\prime}_{\ell}{}^{\top}.$
Thus to conclude the induction step we only need to show that
$\|U^{\prime}_{\ell}\!\Big{(}Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-r_{\ell-1}^{\prime})-Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell})\Big{)}\!U^{\prime}_{\ell}{}^{\top}\|\leq{C_{U}^{2\ell}K\cdot{h}\eta}.$
By the definition of $C_{1}$, the matrices $A=U^{\prime}_{\ell}{}^{\top}U^{\prime}$, $\Delta=-r_{\ell-1}^{\prime}$, satisfy $\max(\|A\|,\|\Delta\|)\leq\|A\|+\|\Delta\|\leq C_{U}^{2}+h^{2}\eta\leq 1+C_{1}h$. Moreover, with the same induction that has led to (28), the polynomial $P_{\ell-1}(x)$ has all its roots bounded by $Lh^{2}C_{\lambda}$, hence $Q_{\ell-1}(x):=P_{\ell-1}(x-b_{\ell-1})$ has all its roots bounded by $(L+1)h^{2}C_{\lambda}\leq 2C_{\lambda}h=C_{0}h$, therefore we can apply Lemma˜K.2 to obtain, with $K=K(C_{0},C_{1})=C_{0}e^{C_{0}(1+C_{1})}+e^{C_{1}}$:
$\|Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-r_{\ell-1}^{\prime})-Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell})\|$ $\leq\frac{K}{h}\|r^{\prime}_{\ell-1}\|\leq K\cdot h\eta$
$\|U^{\prime}_{\ell}\!\Big{(}Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell}-r_{\ell-1}^{\prime})-Q_{\ell-1}(U^{\prime}_{\ell}{}^{\top}U^{\prime}_{\ell})\Big{)}\!U^{\prime}_{\ell}{}^{\top}\|$ $\leq C_{U}^{2}K\cdot h\eta\stackrel{{\scriptstyle C_{U}\geq 1}}{{\leq}}{C_{U}^{2\ell}K\cdot{h}\eta},$
which concludes the induction.
Step 2: Factorisation of $P_{N}$. With the same induction that has led to (28), we have
$P_{N}(x)=\prod_{k=0}^{N-1}\big{(}x-a_{k}\big{)},\qquad a_{0}=0,\quad a_{k}=\sum_{i=1}^{k}h^{2}\lambda_{L-1-i}.$
Applying (34) with $\ell=N$ and recalling $S_{j}=Z_{N}Z_{N}^{\top}$ yields
$S_{j}=P_{N}(M_{N})+E_{N}=\prod_{k=0}^{N-1}\big{(}M_{N}-a_{k}I_{n}\big{)}+E_{N},$
where $\|E_{N}\|\leq C_{U}^{2{N}}NKh\eta\leq(1+C_{1}h)^{{L}}K\eta\leq{\exp(C_{1})}K\eta$.
Since $M_{N}=U^{\prime}_{N}{U^{\prime}_{N}}^{\top}=U_{L-1}{U_{L-1}}^{\top}$, we recover (31) as claimed. ∎
Appendix L Proof of Proposition˜4.7.
See 4.7
Proof.
For convenience we recall the state equation (17) for $Z_{s}$, where $s\in[0,1]$ indicates depth:
$\frac{\mathrm{d}Z_{s}}{\mathrm{d}s}=\mathcal{A}_{s}\,Z_{s},\quad Z_{0}=\mathrm{Id}_{n}\text{ fixed},$ (35)
and we recall that the objective function is factorized by $\ell(\theta)=f\bigl{(}Z_{s=1}\bigr{)}$, where the parameters are the family $\theta=\{\mathcal{A}_{s}:s\in[0,1]\}$.
Let $\theta:[t\in[0,T]\mapsto\theta(t)\in\mathcal{X}]\in\mathcal{C}^{1}([0,T],\mathcal{X})$ be the solution of the gradient flow given by the family of coupled ODE (16)
$\forall s\in[0,1],\quad\frac{\partial\mathcal{A}_{s}}{\partial t}(t)=-\,\mathfrak{g}_{s}(t),\quad\text{with}\quad\mathfrak{g}_{s}(t):=\frac{\partial\ell}{\partial\mathcal{A}_{s}}\Bigl{(}\theta(t)\Bigr{)},$ (36)
with a given initialization $\theta(0)$. Our goal is to show that $\frac{\partial}{\partial t}h_{s}(\theta(t))=0$.
Step 1: Computations of $\frac{\partial}{\partial t}h_{s}(\theta(t))$.
For any $s\in[0,1]$, one has by definition
$h_{s}(\theta(t))=\frac{\partial\mathcal{A}_{s}(t)}{\partial s}+\left(\frac{\partial\mathcal{A}_{s}(t)}{\partial s}\right)^{\top}+\bigl{[}\mathcal{A}_{s}(t)^{\top},\mathcal{A}_{s}(t)\bigr{]}.$
Taking the $t$-derivative yields
$\frac{\partial}{\partial t}h_{s}(\theta(t))$ $=\frac{\partial}{\partial t}\left(\frac{\partial\mathcal{A}_{s}(t)}{\partial s}\right)+\frac{\partial}{\partial t}\left(\frac{\partial\mathcal{A}_{s}(t)}{\partial s}\right)^{\top}+\frac{\partial}{\partial t}\bigl{[}\mathcal{A}_{s}(t)^{\top},\mathcal{A}_{s}(t)\bigr{]}$
$=\frac{\partial}{\partial s}\left(\frac{\partial\mathcal{A}_{s}(t)}{\partial t}\right)+\left(\frac{\partial}{\partial s}\frac{\partial\mathcal{A}_{s}(t)}{\partial t}\right)^{\top}+\frac{\partial}{\partial t}\bigl{[}\mathcal{A}_{s}(t)^{\top},\mathcal{A}_{s}(t)\bigr{]},$ (37)
where the exchange of derivatives is justified in Section˜L.1.
Moreover one has
$\frac{\partial}{\partial t}\bigl{[}\mathcal{A}_{s}(t)^{\top},\mathcal{A}_{s}(t)\bigr{]}=\Bigl{[}\frac{\partial\mathcal{A}_{s}(t)^{\top}}{\partial t},\mathcal{A}_{s}(t)\Bigr{]}+\Bigl{[}\mathcal{A}_{s}(t)^{\top},\frac{\partial\mathcal{A}_{s}(t)}{\partial t}\Bigr{]}.$
Thus by using (36)
$\frac{\partial\mathcal{A}_{s}(t)}{\partial t}=-\,\mathfrak{g}_{s}(t),$
we obtain
$\frac{\partial}{\partial t}h_{s}(\theta(t))$ $=\frac{\partial}{\partial s}\Bigl{(}-\mathfrak{g}_{s}(t)\Bigr{)}+\left(\frac{\partial}{\partial s}\Bigl{(}-\mathfrak{g}_{s}(t)\Bigr{)}\right)^{\top}+\Bigl{[}-\,\mathfrak{g}_{s}(t)^{\top},\mathcal{A}_{s}(t)\Bigr{]}+\Bigl{[}\mathcal{A}_{s}(t)^{\top},-\,\mathfrak{g}_{s}(t)\Bigr{]}$
$=-\frac{\partial\mathfrak{g}_{s}(t)}{\partial s}-\left(\frac{\partial\mathfrak{g}_{s}(t)}{\partial s}\right)^{\top}-\bigl{[}\mathfrak{g}_{s}(t)^{\top},\mathcal{A}_{s}(t)\bigr{]}-\bigl{[}\mathcal{A}_{s}(t)^{\top},\mathfrak{g}_{s}(t)\bigr{]}.$ (38)
The remaining task is to show that the sum of these terms cancels, using an expression of the gradient.
Step 2: An expression of $\mathfrak{g}_{s}(t)$ using the adjoint equation.
To compute the gradient $\mathfrak{g}_{s}=\frac{\partial\ell}{\partial\mathcal{A}_{s}}$, we introduce the adjoint variable (Pontryagin et al. (1962)) $\Lambda_{s}(t)$, which satisfies the adjoint equation
$\frac{\partial\Lambda_{s}(t)}{\partial s}=-\,\mathcal{A}_{s}(t)^{\top}\Lambda_{s}(t),\quad\Lambda_{1}(t)=\frac{\partial f}{\partial Z}\Bigl{(}Z_{1}(t)\Bigr{)}.$ (39)
Moreover it satisfies as shown in Section˜L.2:
$\mathfrak{g}_{s}(t)=\Lambda_{s}(t)\,Z_{s}(t)^{\top}.$ (40)
Step 3: Compute $\frac{\partial}{\partial s}\mathfrak{g}_{s}(t)$.
By differentiating (40) with respect to $s$, we get:
$\frac{\partial}{\partial s}\mathfrak{g}_{s}(t)=\frac{\partial\Lambda_{s}(t)}{\partial s}\,Z_{s}(t)^{\top}+\Lambda_{s}(t)\,\frac{\partial Z_{s}(t)^{\top}}{\partial s}.$
Then by using the adjoint equation (39) and the state equation (35), one has
$\frac{\partial}{\partial s}\mathfrak{g}_{s}(t)$ $=-\,\mathcal{A}_{s}(t)^{\top}\,\Lambda_{s}(t)Z_{s}(t)^{\top}+\Lambda_{s}(t)Z_{s}(t)^{\top}\,\mathcal{A}_{s}(t)^{\top}$ (41)
$=-\,\mathcal{A}_{s}(t)^{\top}\,\mathfrak{g}_{s}(t)+\mathfrak{g}_{s}(t)\,\mathcal{A}_{s}(t)^{\top}$ (42)
$=-[\mathcal{A}_{s}(t)^{\top},\mathfrak{g}_{s}(t)].$ (43)
Taking the transpose,
$\left(\frac{\partial}{\partial s}\mathfrak{g}_{s}(t)\right)^{\top}=-\,\mathfrak{g}_{s}(t)^{\top}\,\mathcal{A}_{s}(t)+\mathcal{A}_{s}(t)\,\mathfrak{g}_{s}(t)^{\top}=[\mathcal{A}_{s}(t),\mathfrak{g}_{s}(t)^{\top}]=-[\mathfrak{g}_{s}(t)^{\top},\mathcal{A}_{s}(t)].$ (44)
Step 4: Conclusion. By substituting the computed expressions into (L), one obtains as claimed that
$\frac{\partial}{\partial t}h_{s}(\theta(t))=0.$
L.1 We now detail (L).
Theorem L.1 (Commutation of mixed derivatives).
Let
$\mathcal{X}\;=\;\mathcal{C}^{1}\!\bigl{(}[0,1],\mathbb{R}^{n\times n}\bigr{)},\qquad\|f\|_{\mathcal{X}}:=\max\!\bigl{\{}\|f\|_{\infty},\,\|f^{\prime}\|_{\infty}\bigr{\}},$
and set $B=\mathcal{C}^{0}([0,1],\mathbb{R}^{n\times n})$ with the sup–norm $\|\cdot\|_{B}=\|\cdot\|_{\infty}$. Denote $D:\mathcal{X}\longrightarrow B,\;f\mapsto f^{\prime}$ the spatial derivative. Suppose $\theta(\cdot)\in\mathcal{C}^{1}\!\bigl{(}[0,T],\mathcal{X}\bigr{)}$ and write $\mathcal{A}(t,s):=[\theta(t)](s)$. Then
  • the mixed derivatives
    $\partial_{t}\partial_{s}\mathcal{A}(t,s)\quad\text{and}\quad\partial_{s}\partial_{t}\mathcal{A}(t,s)$
    exist for every $(t,s)\in[0,T]\times[0,1]$ and coincide:
    $\boxed{\;\partial_{t}\partial_{s}\mathcal{A}(t,s)=\partial_{s}\partial_{t}\mathcal{A}(t,s)\quad\forall\,(t,s)\;}.$
  • the map $s\mapsto\partial_{t}\partial_{s}\mathcal{A}(t,s)$ is continuous.
Proof.
Step 1: $D$ is continuous. For every $f\in\mathcal{X}$,
$\|Df\|_{B}\;=\;\|f^{\prime}\|_{\infty}\;\leq\;\max\!\bigl{\{}\|f\|_{\infty},\|f^{\prime}\|_{\infty}\bigr{\}}=\|f\|_{\mathcal{X}},$
so $\|D\|_{\mathrm{op}}\leq 1$; hence $D$ is a bounded and thus a continuous linear map.
Step 2: Temporal differentiability is preserved by $D$. The fact that the function $\theta$ (valued in the Banach space $\mathcal{X}$) is $\mathcal{C}^{1}$ means precisely that its (Fréchet) derivative $\dot{\theta}(t):=\partial_{t}\theta(t)\in\mathcal{X}$ exists for each $t$ and the map $t\mapsto\dot{\theta}(t)$ is continuous from $[0,T]$ to $\mathcal{X}$.
Applying the continuous and linear operator $D$ yields by linearity
$\frac{D(\theta(t+h))-D(\theta(t))}{h}=D\left(\frac{\theta(t+h)-\theta(t)}{h}\right)$
for every $t\in[0,T]$ and $h$ small enough such that $t+h\in[0,T]$, and since by continuity of $D$ the right hand side tends to $D(\dot{\theta}(t))$ when $h\to 0$, the left hand side also has a limit, showing that
$\frac{d}{dt}\bigl{(}D(\theta(t))\bigr{)}=D\!\bigl{(}\dot{\theta}(t)\bigr{)}\qquad\text{for every }t\in[0,T].$ (45)
Thus the mixed derivative $\partial_{t}\partial_{s}\mathcal{A}(t,\cdot)$ exists as an element of $B$.
Step 3: symmetry of the mixed derivatives. Evaluating the identity (45) above pointwise in $s$ and writing $\mathcal{A}(t,s)=[\theta(t)](s)$ gives
$\partial_{t}\partial_{s}\mathcal{A}(t,s)=\bigl{[}D(\dot{\theta}(t))\bigr{]}(s)=\partial_{s}\bigl{[}\dot{\theta}(t)\bigr{]}(s)=\partial_{s}\partial_{t}\mathcal{A}(t,s).$
Hence the two mixed derivatives exist everywhere and are equal.
Step 4: continuity of $s\mapsto\partial_{t}\partial_{s}\mathcal{A}(t,s)$. Since $\dot{\theta}(t)\in\mathcal{X}$ for each $t$, its derivative $s\mapsto\partial_{s}\bigl{[}\dot{\theta}(t)\bigr{]}(s)$ is continuous. As $\partial_{s}\bigl{[}\dot{\theta}(t)\bigr{]}(s)=\partial_{s}\partial_{t}\mathcal{A}(t,s)$, by the previous step, this exactly means that $s\mapsto\partial_{t}\partial_{s}\mathcal{A}(t,s)$ is continuous. ∎
L.2 We now show (40).
More precisely, to show (40), we will both prove that
$\mathfrak{g}_{s}(t)=(Z_{s}(t)^{-1})^{\top}Z_{1}(t)^{\top}\nabla f(Z_{1}(t))Z_{s}(t)^{\top}$ (46)
and that
$\Lambda_{s}(t)=(Z_{s}(t)^{-1})^{\top}Z_{1}(t)^{\top}\nabla f(Z_{1}(t)),$ (47)
which will indeed give (40).
We briefly explain why, for a given $t$, the matrix $Z_{s}(t)$ never loses its invertibility when $s\in[0,1]$ varies, by showing that the determinant can never reach 0. As
$\partial_{s}Z_{s}(t)=\mathcal{A}_{s}(t)\,Z_{s}(t),\qquad Z_{0}(t)=\mathrm{Id}_{n}.$
Jacobi’s rule gives
$\frac{\mathrm{d}}{\mathrm{d}s}\det Z_{s}(t)=\operatorname{tr}\!\bigl{(}\mathcal{A}_{s}(t)\bigr{)}\,\det Z_{s}(t),\quad\det Z_{0}(t)=1.$
Solving this scalar ODE,
$\det Z_{s}(t)=\exp\!\Bigl{(}\int_{0}^{s}\operatorname{tr}\bigl{(}\mathcal{A}_{\tau}(t)\bigr{)}\,d\tau\Bigr{)}\neq 0,\qquad s\in[0,1].$
Therefore $Z_{s}(t)\in\mathrm{GL}(n)$ for every $s$.
Since $t$ is fixed, in the following we lighten notations by dropping it from the equations. The proof of (47) is direct by showing that $\Lambda_{s}$ and $(Z_{s}^{-1})^{\top}Z_{1}^{\top}\nabla f(Z_{1})$ satisfy the same ODE (39) with the same value at $s=1$. Thus we only need to show (46).
Proof.
To show (46) we will use Riesz theorem to identify the expression of the gradient. We thus will consider the Hilbert space
$L^{2}:=L^{2}\bigl{(}[0,1],\mathbb{R}^{n\times n}\bigr{)},\qquad\langle U,V\rangle_{L^{2}}:=\int_{0}^{1}\!\operatorname{tr}\!\bigl{(}U_{s}^{\top}V_{s}\bigr{)}\,ds,$
in which the parameter $\theta=\{\mathcal{A}_{s}\in\mathbb{R}^{n\times n}:s\in[0,1]\}\in\mathcal{C}^{1}([0,1],\mathbb{R}^{n\times n})=:\mathcal{X}\subseteq L^{2}$ lives.
We recall that $Z_{s}(\theta)$ is the unique solution of the state equation (17):
$\partial_{s}Z_{s}=\mathcal{A}_{s}Z_{s},\quad Z_{0}=\mathrm{Id}_{n},\qquad\forall s\in[0,1],$ (48)
and that the cost $\ell$ is factorized by the flow map $Z_{1}(\theta)$ with a smooth scalar field $f:\mathbb{R}^{n\times n}\to\mathbb{R}$, i.e, $\ell(\theta):=f\!\bigl{(}Z_{1}(\theta)\bigr{)}$.
1st step: expression of the Gateaux variation of the flow
Let $\theta=\{\mathcal{A}_{s}:s\in[0,1]\}\in\mathcal{X}$ be fixed and pick an arbitrary $\delta\theta\in\mathcal{X}$. For $\varepsilon\in\mathbb{R}$ define the perturbed coefficient $\theta^{\varepsilon}:=\theta+\varepsilon\,\delta\theta$, denoting its components $\theta^{\varepsilon}=\{\mathcal{A}_{s}^{\epsilon}:s\in[0,1]\}$. Denote by $Z_{s}^{\varepsilon}:=Z_{s}(\theta^{\varepsilon})$ the flow that satisfies the associated ODE:
$\partial_{s}Z_{s}^{\varepsilon}=\mathcal{A}_{s}^{\epsilon}Z_{s}^{\epsilon},\,Z_{0}^{\epsilon}=\mathrm{Id}_{n},\qquad\forall s\in[0,1].$ (49)
As $(s,\epsilon,Z)\mapsto\mathcal{A}_{s}^{\epsilon}Z\in\mathcal{C}^{1}$, th function $(s,\epsilon)\mapsto Z_{s}^{\epsilon}$ is $\mathcal{C}^{1}$ using the Cauchy–Lipschitz theorem with a parameter. In particular for any $s\in[0,1]$, $\epsilon\mapsto Z_{s}^{\epsilon}$ is $\mathcal{C}^{1}$. Introduce the first variation
$\delta Z_{s}=\left.\frac{d}{d\varepsilon}Z_{s}^{\varepsilon}\right|_{\varepsilon=0}=:\Delta_{s},$
which corresponds to the Gateaux derivative of $\theta^{\prime}\mapsto Z_{s}(\theta^{\prime})$ at $\theta$ in the direction $h=\delta\theta$. We now show that $\Delta_{s}$ satisfies the following inhomogeneous ODE:
$\partial_{s}\Delta_{s}=\mathcal{A}_{s}\Delta_{s}+\delta\mathcal{A}_{s}\,Z_{s},\qquad\Delta_{0}=0,\qquad\forall s\in[0,1].$ (50)
where $\delta\mathcal{A}_{s}\coloneqq\left.\frac{d}{d\varepsilon}\mathcal{A}_{s}^{\varepsilon}\right|_{\varepsilon=0}$.
Indeed let us consider $q_{s}^{\varepsilon}\coloneqq\frac{Z_{s}^{\varepsilon}-Z_{s}}{\varepsilon}$ for any $0<\varepsilon\leq 1$. In particular one has $q_{s}^{\varepsilon}\underset{\varepsilon\rightarrow 0}{\longrightarrow}\Delta_{s}$. Moreover one has:
$q_{s}^{\varepsilon}=\varepsilon^{-1}\int_{0}^{s}(\mathcal{A}_{u}^{\varepsilon}Z_{u}^{\varepsilon}-\mathcal{A}_{u}Z_{u})\mathrm{d}u=\int_{0}^{s}B_{u}^{\epsilon}Z_{u}\mathrm{d}u+\int_{0}^{s}\mathcal{A}_{u}^{\varepsilon}q_{u}^{\varepsilon}\mathrm{d}u,$ (51)
where $B_{u}^{\epsilon}\coloneqq\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}}{\varepsilon}=\frac{\mathcal{A}_{u}^{\varepsilon}-\mathcal{A}_{u}^{0}}{\varepsilon}$ satisfies $B_{u}^{\varepsilon}\underset{\varepsilon\rightarrow 0}{\longrightarrow}\left.\frac{d}{d\varepsilon}\mathcal{A}_{u}^{\varepsilon}\right|_{\varepsilon=0}=\delta\mathcal{A}_{u}$ (as $\epsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$) and where $\varepsilon\in[0,1]\mapsto B_{u}^{\varepsilon}$ is continuous (at 0, we define $B_{u}^{0}=\delta\mathcal{A}_{u}$) as $\varepsilon\mapsto\mathcal{A}_{u}^{\varepsilon}$ is $\mathcal{C}^{1}$), and thus is bounded on $[0,1]$ by a constant that does not depend on $\varepsilon$. By dominated convergence, when $\varepsilon\rightarrow 0$ in (51) one obtains the limit:
$\Delta_{s}=\int_{0}^{s}(\mathcal{A}_{u}\Delta_{u}+\delta\mathcal{A}_{u}\,Z_{u})\mathrm{d}u,$
which coincides with the unique solution of (50).
Since $Z_{s}$ is a solution for the homogeneous part $\partial_{s}Z_{s}=\mathcal{A}_{s}Z_{s}$ with $Z_{0}=\mathrm{Id}_{n}$, by the variation-of-parameters method, one obtains (as $\Delta_{0}=0$):
$\Delta_{s}=Z_{s}\int_{0}^{s}Z_{\tau}^{-1}\,\delta\mathcal{A}_{\tau}\,Z_{\tau}\,d\tau.$
Evaluating at $s=1$ gives
$\delta Z_{1}=\Delta_{1}=\int_{0}^{1}Z_{1}Z_{\tau}^{-1}\,\delta\mathcal{A}_{\tau}\,Z_{\tau}\,d\tau.$ (52)
2d step: Differential of $\ell$ and identification of the gradient.
Because $f\in\mathcal{C}^{1}(\mathbb{R}^{n\times n},\mathbb{R})$ its (Fréchet) differential at $M\in\mathbb{R}^{n\times n}$ is
$Df(M)[H]=\langle\nabla f(M),H\rangle_{F},\qquad\forall\,H\in\mathbb{R}^{n\times n}.$ (53)
Applying the chain rule to $\ell=f\circ Z_{1}$ with the Gateaux differentials $D_{G}$ at $\theta$ and in the direction $h=\delta\theta$, one obtains,
$D_{G}\ell(\theta)[\delta\theta]=D_{G}f\bigl{(}Z_{1}(\theta)\bigr{)}\bigl{[}\,\delta Z_{1}\,\bigr{]}.$
But as by hypothesis both $\ell$ and $f$ are Frechet differentiable, one has:
$D\ell(\theta)[\delta\theta]=Df\bigl{(}Z_{1}(\theta)\bigr{)}\bigl{[}\,\delta Z_{1}\,\bigr{]}.$
Using (53) with $M=Z_{1}(\theta)$ and $H=\delta Z_{1}$,
$D\ell(\theta)[\delta\theta]=\bigl{\langle}\nabla f\!\bigl{(}Z_{1}(\theta)\bigr{)},\delta Z_{1}\bigr{\rangle}_{F}.$ (54)
By inserting the expression of $\delta Z_{1}$ from (52) into (54), one has:
$D\ell(\theta)[\delta\theta]=\int_{0}^{1}\operatorname{tr}\!\Bigl{(}\nabla f(Z_{1})^{\top}\,Z_{1}Z_{\tau}^{-1}\,\delta\mathcal{A}_{\tau}\,Z_{\tau}\Bigr{)}\,d\tau.$
Because $\operatorname{tr}(RS)=\operatorname{tr}(SR)$, one get
$\operatorname{tr}\!\bigl{(}\nabla f(Z_{1})^{\top}Z_{1}Z_{\tau}^{-1}\delta\mathcal{A}_{\tau}Z_{\tau}\bigr{)}=\operatorname{tr}\!\bigl{(}Z_{\tau}\,\nabla f(Z_{1})^{\top}Z_{1}Z_{\tau}^{-1}\,\delta\mathcal{A}_{\tau}\bigr{)},$
and thus by defining for each $\tau$
$G_{\tau}^{\top}:=Z_{\tau}\,\nabla f(Z_{1})^{\top}Z_{1}Z_{\tau}^{-1},$ (55)
one finally has
$D\ell(\theta)[\delta\theta]=\int_{0}^{1}\operatorname{tr}\bigl{(}G_{\tau}^{\top}\,\delta A_{\tau}\bigr{)}\,d\tau=\langle G,\delta\theta\rangle_{L^{2}}.$
By Riesz theorem, the gradient in $L^{2}$ (i.e. the Fréchet gradient) is the unique element $G\in L^{2}$ verifying $D\ell(\theta)[\delta\theta]=\langle G,\delta\theta\rangle_{L^{2}}$ for every $\delta\theta$:
$\nabla\ell(\theta)=G.$
The transpose in (55) finally yields the required formula. ∎
L.3 Link with conservation laws in finite depth (informal).
We assume that $\theta_{L}\coloneqq(U_{L},\cdots,U_{1})$ satisfies the relaxed balanced conditions:
$U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}=\frac{H_{i}}{L^{2}},$
then as $U_{k}=\mathrm{Id}+\frac{1}{L}A_{k}$, using that $A_{k+1}=A_{k}+\frac{1}{L}A_{k}^{\prime}+o\left(\frac{1}{L}\right)$ we get:
$\frac{H_{k}}{L^{2}}$ $=(\mathrm{Id}+\frac{1}{L}A_{k+1})^{\top}(\mathrm{Id}+\frac{1}{L}A_{k+1})-(\mathrm{Id}+\frac{1}{L}A_{k})(\mathrm{Id}+\frac{1}{L}A_{k}^{\top})$
$=\frac{A_{k+1}^{\top}+A_{k+1}}{L}-\frac{A_{k}+A_{k}^{\top}}{L}-\frac{A_{k}A_{k}^{\top}}{L^{2}}+\frac{A_{k+1}^{\top}A_{k+1}}{L^{2}}$
$=\frac{A_{k}^{\prime\top}+A_{k}^{\prime}-A_{k}A_{k}^{\top}+A_{k}^{\top}A_{k}}{L^{2}}+o\left(\frac{1}{L^{2}}\right)$
$=\frac{h_{s_{k}}(\theta)}{L^{2}}+o\left(\frac{1}{L^{2}}\right),$
Thus $h_{s}$ is such that $h_{s_{k}}(\theta)=H_{k}+o(1)$.
In particular if $\theta_{L}$ satisfies the relaxed balanced condition
$U_{i+1}^{\top}U_{i+1}-U_{i}U_{i}^{\top}=\frac{\lambda_{i}}{L^{2}}\mathrm{Id},$
then one can choose $h_{s}$ as:
$h_{s}(\theta)=\lambda(s)\mathrm{Id}_{n},$
with $\lambda$ a function such that $\lambda(s_{k})=\lambda_{k}$. We say in that case that $\theta$ satisfies the relaxed balanced condition.
Appendix M Proof of Theorem˜4.8.
M.1 Proof of the theorem
See 4.8
Proof.
For any $t$ and any integer $L\geq 1$ we define $s_{k}:=s_{k}^{L}=\tfrac{k}{L}$ for $k=0,\cdots,L-1$ and:
$X_{k+1}(t)=X_{k}(t)+h\mathcal{A}_{s_{k}}(t)X_{k}(t),\quad{\text{with}\ h:=\frac{1}{L}\ \text{and}}\ X_{0}(t)=I_{n}.$
Since this corresponds exactly to the Euler explicit method with step $h$ for the ODE
$\partial_{s}Z_{s}(t)=\mathcal{A}_{s}(t)Z_{s}(t),\quad Z_{0}(t)=\mathrm{Id}_{n},\quad s\in[0,1],$
one has for any $t$ and $L$ (computations postponed in Section˜M.2):
$\sup_{0\leq k\leq L-1}\|X_{k}(t)-Z_{s_{k}}(t)\|$ $=\mathcal{O}(h)$ (56)
$\|\partial_{t}Z_{1}(t)-\partial_{t}X_{L}(t)\|$ $=\mathcal{O}(h),$ (57)
with $\|\cdot\|$ any matrix norm on $\mathbb{R}^{n\times n}$, the implicit constant in the notation $\mathcal{O}(\cdot)$ is independent of $k$ and $L$ while it can depend on $t$.
We now fix some $t$ and observe that $X_{k+1}(t)=U_{k}(t)X_{k}(t)$ with
$U_{k}(t)\coloneqq\mathrm{Id}_{n}+h\mathcal{A}_{s_{k}}(t).$ (58)
so that (from now on we drop the $t$ variable for brevity)
$\partial_{t}X_{L}=h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(\partial_{t}\mathcal{A}_{s_{j}})X_{j}(t)=h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(\partial_{t}\mathcal{A}_{s_{j}})U_{j-1}\cdots U_{0}$
By (16) and the relation (46) (shown in Section˜L.2) we have for any $s\in[0,1]$
$\partial_{t}\mathcal{A}_{s}=-\mathfrak{g}_{s}=-(Z_{s}^{-1})^{\top}Z_{1}^{\top}\nabla f(Z_{1})Z_{s}^{\top}$
hence
$\partial_{t}X_{L}=-h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(Z_{s_{j}}^{-1})^{\top}Z_{1}^{\top}\nabla f(Z_{1})Z_{s_{j}}^{\top}U_{j-1}\cdots U_{0}$
As $U_{L-1}\cdots U_{j+1}=X_{L}X_{j+1}^{-1}=(Z_{1}+\mathcal{O}(h))(Z_{s_{j+1}}+\mathcal{O}(h))^{-1}=Z_{1}Z_{s_{j+1}}^{-1}+\mathcal{O}(h)$ since the invertibility of $Z_{s}$ and continuity of $s\mapsto Z_{s}$ implies that $\|Z_{s}^{-1}\|$ is uniformly bounded) and $Z_{s_{j+1}}Z_{s_{j}}^{-1}=\mathrm{Id}_{n}+\mathcal{O}(h)$ (since $Z_{s_{j+1}}=Z_{s_{j}}+h\mathcal{A}_{s_{j}}Z_{s_{j}}+\mathcal{O}(h^{2})$), we deduce that
$(Z_{s_{j}}^{-1})^{\top}Z_{1}^{\top}$ $=(Z_{1}Z_{s_{j}}^{-1})^{\top}=[(Z_{1}Z_{s_{j+1}}^{-1})Z_{s_{j+1}}Z_{s_{j}}^{-1}]^{\top}$
$=[(U_{L-1}\cdots U_{j+1}+\mathcal{O}(h))(I_{n}+\mathcal{O}(h))]^{\top}$
$=(U_{L-1}\cdots U_{j+1})^{\top}+\mathcal{O}(h),$
where in the last line we used that since with any relevant matrix norm since $\max_{k}\|U_{k}\|=1+\mathcal{O}(h)=1+\mathcal{O}(1/L)$ we have $\|U_{L-1}\ldots U_{j}\|\leq[1+\mathcal{O}(1/L)]^{L}=\mathcal{O}(1)$. Similarly we also have $\|U_{j-1}\ldots U_{0}\|=\mathcal{O}(1)$ hence
$\partial_{t}X_{L}=$ $-h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(U_{L-1}\cdots U_{j+1})^{\top}\nabla f(Z_{1})Z_{s_{j}}^{\top}U_{j-1}\cdots U_{0}+\underbrace{h\sum_{j=0}^{L-1}\mathcal{O}(h)}_{=\mathcal{O}(h)\ \text{since}\ h=1/L}$
Similarly as $U_{j-1}\cdots U_{0}=X_{j}=Z_{s_{j}}+\mathcal{O}(h)$ by (56), we get $Z_{s_{j}}^{\top}=(U_{j-1}\cdots U_{0})^{\top}+\mathcal{O}(h)$ hence
$\partial_{t}X_{L}=$ $-h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(U_{L-1}\cdots U_{j+1})^{\top}\nabla f(Z_{1})(U_{j-1}\cdots U_{0}+\mathcal{O}(h))^{\top}(U_{j-1}\cdots U_{0})$
$=$ $-h\sum_{j=0}^{L-1}(U_{L-1}\cdots U_{j+1})(U_{L-1}\cdots U_{j+1})^{\top}\nabla f(Z_{1})(U_{j-1}\cdots U_{0})^{\top}(U_{j-1}\cdots U_{0})+\mathcal{O}(h)$ (59)
We also have
$U_{k+1}^{\top}U_{k+1}-U_{k}U_{k}^{\top}$ $=(\mathrm{Id}_{n}+h\mathcal{A}_{s_{k+1}})^{\top}(\mathrm{Id}_{n}+h\mathcal{A}_{s_{k+1}})-(\mathrm{Id}_{n}+h\mathcal{A}_{s_{k}})(\mathrm{Id}_{n}+h\mathcal{A}_{s_{k}}^{\top})$
$=h(\mathcal{A}_{s_{k+1}}^{\top}+\mathcal{A}_{s_{k+1}})-h(\mathcal{A}_{s_{k}}+\mathcal{A}_{s_{k}}^{\top})-h^{2}(\mathcal{A}_{s_{k}}\mathcal{A}_{s_{k}}^{\top})+h^{2}(\mathcal{A}_{s_{k+1}}^{\top}\mathcal{A}_{s_{k+1}})$
$=h^{2}(\mathcal{A}_{s_{k}}^{\prime\top}+\mathcal{A}_{s_{k}}^{\prime}-\mathcal{A}_{s_{k}}\mathcal{A}_{s_{k}}^{\top}+\mathcal{A}_{s_{k}}^{\top}\mathcal{A}_{s_{k}})+{o}\left(h^{2}\right)$
$=h^{2}\mathbf{h}_{s_{k}}(\theta)+{o}\left(h^{2}\right)$
$=h^{2}\lambda(s_{k})\mathrm{Id}_{n}+{o}\left(h^{2}\right),$ (60)
as $\theta(0)$ satisfies the quasi-balanced condition and $\mathcal{A}_{s_{k+1}}=\mathcal{A}_{s_{k}}+h\mathcal{A}_{s_{k}}^{\prime}+{o}\left(h\right)$, where the implicit $o(1)$ function in the notation $o(h^{2})=h^{2}o(1)$ is still independent of $k$ and $L$ as $s\in[0,1]\mapsto\mathcal{A}_{s}^{\prime}$ is continuous on the compact set $[0,1]$ and is thus uniformly continuous.
By Lemma˜K.1 with $\lambda_{k}=\lambda(s_{k})$ and one has:
$(U_{L-1}\cdots U_{j+1})(U_{L-1}\cdots U_{j+1})^{\top}=\prod_{k=0}^{L-1-(j+1)}(U_{L-1}U_{L-1}^{\top}-a_{k}\mathrm{Id}_{n})+E_{L-1-j},$ (61)
with
$a_{0}\coloneqq 0,\quad a_{k}=h^{2}\sum_{i=1}^{k}\lambda(s_{L-1}-s_{i})\ \text{for}\ k\geq 1,\quad\text{and}\ \|E_{L-1-j}\|\leq K\eta$ (62)
where $K\coloneqq(C_{0}\exp(C_{1})\exp(C_{0}(C_{1}+1)+\exp(C_{1}))$ with $C_{0}\coloneqq 2C_{\lambda}$, $C_{1}\coloneqq(C_{U}^{2}+\eta h^{2}-1)/h$,
$C_{U}$ $\coloneqq\max(1,\max_{k}\|U_{k}\|),\quad C_{\lambda}\coloneqq\max_{k}|\lambda_{k}|$ (63)
$\eta$ $\coloneqq L^{2}\cdot\max_{0\leq k\leq L-2}\|(U_{k+1}^{\top}U_{k+1}-U_{k}U_{k}^{\top})-h^{2}\lambda_{k}\,\mathrm{Id}_{n}\|.$ (64)
As $\lambda(\cdot)$ is continuous, $C_{\lambda}\leq\|\lambda\|_{\infty}<\infty$ for any $L$. Similarly, we already used that as $s\in[0,1]\mapsto\mathcal{A}_{s}$ is continuous, $C_{U}=1+\mathcal{O}(h)$, and thus $C_{U}^{2}=1+\mathcal{O}(h)$, again with implicit constant independent of $L$. Moreover by (M.1), $\eta h^{2}=\eta/L^{2}=o(h^{2})$, and we obtain $C_{1}=(C_{U}^{2}+\eta h^{2}-1)/h=(\mathcal{O}(h)+o(h^{2}))/h=\mathcal{O}(1)$, hence $C_{1}$ is bounded uniformly. Finally we obtain
$\max_{j}E_{L-1-j}=o(1)$ (65)
where the implicit function $o(1)$ is still independent of $L$.
We denote
$F_{i}(U_{L-1}U_{L-1}^{\top})\coloneqq\prod_{k=0}^{i}(U_{L-1}U_{L-1}^{\top}-a_{k}\mathrm{Id}_{n})$ (66)
and use the shorthand $A_{k}:=A_{k}^{L}(t)\coloneqq\mathcal{A}_{s_{k}}(t)\in\mathbb{R}^{n\times n}$, for $0\leq k\leq L-1$. Since $U_{k}=\mathrm{Id}_{n}+hA_{k}$ with $h=1/L$ and $s_{L-1}-s_{i}=\frac{L-1}{L}-\frac{i}{L}=1-\frac{i+1}{L}$ for each integer $i$, we have (using Riemann integration as $\lambda(\cdot)$ is continuous): since all the matrices in the product (66) commute
$F_{j}(U_{L-1}U_{L-1}^{\top})$ $=\exp\left(\sum_{k=0}^{j}\log(U_{L{-1}}U_{L{-1}}^{\top}-a_{k}\mathrm{Id}_{n})\right)$
[…]
$=\exp\left(\sum_{k=0}^{j}\left(\frac{A_{L-1}+A_{L-1}^{\top}}{L}-\frac{1}{L}\int_{0}^{s_{k}}\lambda(1-v)\mathrm{d}v\cdot\mathrm{Id}_{n}+o\left(\frac{1}{L}\right)\right)\right)$
$=\exp\left(s_{j}(\mathcal{A}_{1}+\mathcal{A}_{1}^{\top})-\underbrace{\frac{1}{L}\sum_{k=0}^{j}\int_{0}^{s_{k}}\lambda(1-v)\mathrm{d}v}_{=\int_{0}^{s_{j}}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u{+o(1)}}\cdot\mathrm{Id}_{n}+o\left(1\right)\right).$
We denote $\psi_{1}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(1-v)\mathrm{d}v\mathrm{d}u$. By (61) and the above derivations one has
$Z_{1}Z_{1}^{\top}=\underset{L\rightarrow+\infty}{\lim}X_{L}X_{L}^{\top}=\underset{L\rightarrow+\infty}{\lim}F_{L-1}(U_{L-1}U_{L-1}^{\top})=\exp((\mathcal{A}_{1}+\mathcal{A}_{1}^{\top})-\psi_{1}(1)\cdot\mathrm{Id}_{n}),$
and thus
$\mathcal{A}_{1}+\mathcal{A}_{1}^{\top}=\log(Z_{1}Z_{1}^{\top})+\psi_{1}(1)\cdot\mathrm{Id}_{n}$
Thus
$F_{j}(U_{L-1}U_{L-1}^{\top})=(Z_{1}Z_{1}^{\top})^{s_{j}}\exp\left(s_{j}\psi_{1}(1)-\psi_{1}(s_{j}){+o(1)}\right).$ (67)
Similarly as before and by adapting the proof of Lemma˜K.1 one gets:
$(U_{j-1}\cdots U_{0})^{\top}(U_{j-1}\cdots U_{0})=\prod_{k=0}^{j-1}(U_{0}^{\top}U_{0}-b_{k}\mathrm{Id}_{n})+o(1)=:G_{j}(U_{0}^{\top}U_{0})+o(1)$ (68)
with $b_{k}=-h^{2}\sum_{i=0}^{k-1}\lambda(s_{i})$ and $b_{0}=0$, and where we denote
$G_{j}(U_{0}^{\top}U_{0})\coloneqq\prod_{k=0}^{j-1}(U_{0}^{\top}U_{0}-b_{k}\mathrm{Id}_{n})$ (69)
Similarly as above:
$G_{j+1}(U_{0}^{\top}U_{0})$ $=\exp\left(\sum_{k=0}^{j}\log(U_{0}^{\top}U_{0}-b_{k}\mathrm{Id}_{n})\right)$
$=\exp\left(\sum_{k=0}^{j}\log\left(\mathrm{Id}_{n}+\frac{A_{0}+A_{0}^{\top}}{L}+o\left(\frac{1}{L}\right)+\frac{1}{L}\left(\underbrace{\frac{1}{L}\sum_{i=0}^{k-1}\lambda(s_{i})}_{=\int_{0}^{s_{k}}\lambda(v)\mathrm{d}v+o(1)}\right)\mathrm{Id}_{n}\right)\right)$
$=\exp\left(\sum_{k=0}^{j}\left(\frac{A_{0}+A_{0}^{\top}}{L}+\frac{1}{L}\int_{0}^{s_{k}}\lambda(v)\mathrm{d}v\cdot\mathrm{Id}_{n}+o\left(\frac{1}{L}\right)\right)\right)$
$=\exp\left(s_{j}(\mathcal{A}_{0}+\mathcal{A}_{0}^{\top})+\underbrace{\frac{1}{L}\sum_{k=0}^{j}\int_{0}^{s_{k}}\lambda(v)\mathrm{d}v}_{=\int_{0}^{s_{j}}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u{+o(1)}}\cdot\mathrm{Id}_{n}+o(1)\right).$
We denote $\psi_{2}:s\in[0,1]\mapsto\int_{0}^{s}\int_{0}^{u}\lambda(v)\mathrm{d}v\mathrm{d}u$. By (68) and the above derivations we have
$Z_{1}^{\top}Z_{1}=\underset{L\rightarrow+\infty}{\lim}X_{L}^{\top}X_{L}=\underset{L\rightarrow+\infty}{\lim}G_{L}(U_{0}^{\top}U_{0})=\exp((\mathcal{A}_{0}+\mathcal{A}_{0}^{\top})+\psi_{2}(1)\cdot\mathrm{Id}_{n}),$
and thus
$\mathcal{A}_{0}+\mathcal{A}_{0}^{\top}=\log(Z_{1}^{\top}Z_{1}){-}\psi_{2}(1)\cdot\mathrm{Id}_{n}$
It follows that
$G_{j+1}(U_{0}^{\top}U_{0})=(Z_{1}^{\top}Z_{1})^{s_{j}}\exp(-{s_{j}}\psi_{2}(1)+\psi_{2}(s_{j}){+o(1)}).$ (70)
Finally, combining (61)-(65)-(66)-(67) and (68)-(69)-(70), we obtain
[…]
$=-h\sum_{j=0}^{L-1}\left[(Z_{1}Z_{1}^{\top})^{1-s_{j}}\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s_{j}}\cdot\exp(\gamma(s))+o(1)\right]+\mathcal{O}(h)$
$=-\int_{0}^{1}(Z_{1}Z_{1}^{\top})^{1-s}\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s}\exp(\gamma(s))\mathrm{d}s+o(1),$
where
$\gamma(s)\coloneqq(1-s)\psi_{1}(1)-\psi_{1}(1-s)-s\psi_{2}(1)+\psi_{2}(s).$
Thus one has:
$\partial_{t}Z_{1}(t)={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-}\int_{0}^{1}(Z_{1}Z_{1}^{\top})^{1-s}\exp(\gamma(s))\nabla f(Z_{1})(Z_{1}^{\top}Z_{1})^{s}\mathrm{d}s,$
which concludes the proof. ∎
M.2 Proof of (56)-(57)
We now show that  (56)-(57) hold for any $t$.
Proof.
First we recall that
$X_{k+1}(t)=X_{k}(t)+h\mathcal{A}_{s_{k}}(t)X_{k}(t),$
with $X_{0}(t)=\mathrm{Id}_{n}$. This corresponds exactly to the Euler explicit formulation of the ODE:
$\partial_{s}Z_{s}(t)=\mathcal{A}_{s}(t)Z_{s}(t),\quad Z_{0}(t)=\mathrm{Id}_{n},\quad s\in[0,1]$ (71)
with step $h=1/L$.
We now show both items at once. We fix some $t$. Set
$W(s)\;:=\;\begin{pmatrix}Z_{s}\\ Y_{s}\end{pmatrix}\in\mathbb{R}^{2n\times n},\qquad Y_{s}:=\partial_{t}Z_{s},\qquad W(0)=\begin{pmatrix}\mathrm{Id}_{n}\\ 0\end{pmatrix}$,
so that
$\frac{\mathrm{d}}{\mathrm{d}s}W(s)=\begin{pmatrix}\mathcal{A}_{s}(t)&0\\[4.0pt] \partial_{t}\mathcal{A}_{s}(t)&\mathcal{A}_{s}(t)\end{pmatrix}\!W(s).$ (72)
The corresponding explicit-Euler discretization with step $h=1/L$ reads
$W_{k+1}=W_{k}+h\,\begin{pmatrix}\mathcal{A}_{s_{k}}(t)&0\\[4.0pt] \partial_{t}\mathcal{A}_{s_{k}}(t)&\mathcal{A}_{s_{k}}(t)\end{pmatrix}W_{k},$ (73)
which coincides component-wise with the recursions for $X_{k}$ and ${T_{k}}=\partial_{t}X_{k}$.
Because the right-hand side of (72) $(s,W)\mapsto\begin{pmatrix}\mathcal{A}_{s}(t)&0\\[4.0pt] \partial_{t}\mathcal{A}_{s}(t)&\mathcal{A}_{s}(t)\end{pmatrix}\!W$ is $\mathcal{C}^{1}$ (indeed both $s\mapsto\mathcal{A}_{s}(t)$ and $s\mapsto\partial_{t}\mathcal{A}_{s}(t)$ are $C^{1}$ (cf Theorem˜L.1) for each $t$), the Euler explicit scheme converges at order one (see e.g. (Berthelin, 2017, Proposition 10.30)):
$\max_{0\leq k\leq L}\,\bigl{\|}W(s_{k})-W_{k}\bigr{\|}\;=\;\mathcal{O}(h).$
In particular one get that for any $k$: $X_{k}(t)=Z_{s_{k}}(t)+O(h)$ and (reading the second bloc row at the final index $k=L$):
$\bigl{\|}\partial_{t}Z_{{s=}1}(t)-\partial_{t}X_{L}(t)\bigr{\|}{=\|Y_{s=1}-T_{L}\|}=\mathcal{O}(h).$